### v2

Abstract.
In this paper, we provide a criterion for the effectiveness and ampleness of divisors on fake quadrics. As applications, first we prove that the fake quadric $S$ satisfies the bounded cohomology property, which is that there exists a positive constant $c_{S}$ such that
$h^{1}(\mathcal{O}_{S}(C))\leq c_{S}h^{0}(\mathcal{O}_{S}(C))$
for every curve $C$ on $S$. And then we give that any fake quadric cannot be embedded in $\mathbb{P}^{4}$. Ultimately, we get that any simply-connected fake quadric $S$ is not a Mori dream space, and the geometric genus $p_{g}(C)>2$ for any curve $C\subset S$.
1. introduction
The study of fake quadrics is an important topic of algebraic surfaces. A fake quadric $S$ is a minimal projective surface that has the same Betti numbers as $\mathbb{P}^{1}\times\mathbb{P}^{1}$ but is not biholomorphic to one, or equivalently a minimal surface of general type with numerical invariants:
$q(S)=p_{g}(S)=0,K_{S}^{2}=8,$
where $q(S)$ is the irregularity, $p_{g}(S)$ is the geometric genus and $K_{S}$ is the canonical divisor of $S$.
Fake quadrics are motivated by the study of fake projective planes, that is, surfaces of general type with the same Betti numbers as $\mathbb{P}^{2}$. It is proved by Yau that every fake projective plane is a quotient $\mathbb{B}_{2}/\Gamma$ of the two-dimensional complex ball $\mathbb{B}_{2}$ by a cocompact and torsion-free discrete group $\Gamma$ of Aut$(\mathbb{B}_{2})$ (c.f.[30],[31]). Based on the Yau’s Theorem, a classification of all fake projective planes is done by Prasad and Yeung (c.f.[26]). However, the classification of fake quadrics seems far from being reached.
Let $S$ be a fake quadric. It is well-known that the canonical divisor $K_{S}$ is ample. In fact, for any reduced and irreducible curve $C$ on $S$, since $K_{S}$ is nef, so $K_{S}.C=0$ if and only if $C$ is a $(-2)$-curve. Hence, $K_{S}$ is ample if and only if there is no $(-2)$-curve exists. Since the number $k$ of $(-2)$-curve is at most $\rho(S)-1=1$ (c.f. [4, VII. Proposition 2.5]), and $k\leq 2(3c_{2}-c_{1}^{2})/9=4/9<1$ (c.f [24, Proposition 2.1.1]), thus $k=0$ which implies $K_{S}$ is ample.
Let $\mathrm{Pic}^{n}(S)$ be the subgroup of $\mathrm{Pic}(S)$ of divisor classes numerically equivalent to zero, denote by $D\equiv 0$ (i.e., $D\in\mathrm{Pic}^{n}(S)$ if $D.E=0$ for all divisors $E$). Let $\mathrm{N}^{1}(S):=\mathrm{Pic}(S)/\mathrm{Pic}^{n}(S)$ be the Néron–Severi lattice of $S$. By the exponential exact sequence, the first Chern class induces an isomorphism $\mathrm{Pic}(S)\cong H^{2}(S,\mathbb{Z})$. Hence the Néron–Severi lattice $\mathrm{N}^{1}(S)\cong H^{2}(S,\mathbb{Z})/\mathrm{tor}$. Furthermore, the Poincaré duality and Hodge index theorem imply that $\mathrm{N}^{1}(S)$ is a unimodular indefinite rank $2$ lattice of signature 00. As a consequence, the Néron–Severi lattice must be even or odd, which means there exist only two cases:
$\mathrm{N}^{1}(S)=U:=\left(\begin{array}[]{cc}0&1\\ 1&0\\ \end{array}\right)\;\;\textrm{ or }\;\;\mathrm{N}^{1}(S)=\langle 1\rangle\oplus\langle-1\rangle:=\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$
For convenience, by a fake quadric of even type we mean its Néron–Severi lattice is an even lattice (i.e., the hyperbolic plane $U$), and by a fake quadric of odd type we mean its Néron–Severi lattice is an odd lattice (i.e., $\langle 1\rangle\oplus\langle-1\rangle$).
To further understand the fake quadrics and clarify their properties, in this paper we mainly focus on the properties of divisors on such surfaces. Specifically,
  1. (1)
    we establish a criterion for evaluating the effectiveness of divisors;
  2. (2)
    we discuss the sufficient and necessary conditions for determining the ample divisors.
We will study these two problems for fake quadrics of even type in Theorem 2.2 of Sec.2 and for fake quadrics of odd type in Theorem 3.7 of Sec.3. The effectiveness and ampleness of divisors on quadrics have never been precisely discussed until now. We believe that this result is new, particularly in those of odd type.
In previous work on the fake quadrics, the odd type was rarely involved. We believe that the possibility of the odd type cannot be ruled out by using the current methods and tools. This led us to have a detailed discussion on this issue. In particular, we provide a detailed argument for the existence of negative curves of the odd type in Sect.3.
Recall that a smooth projective surface $S$ satisfies bounded cohomology property if there is a positive constant $c_{S}$ such that
$h^{1}(\mathcal{O}_{S}(C))\leq c_{S}h^{0}(\mathcal{O}_{S}(C))$
for every curve $C$ on $S$ (Here the curve $C$ is reduced and irreducible). It arose out of discussions between Harbourne, Roé, Cilberto and Miranda (cf.[3]). In [3], the authors disproved such property by giving a counterexample of surfaces of general type with a large Picard number (cf. [2, Corollary 3.1.2]). However, it could still be true for rational surfaces (cf.[3]) and some special projective surfaces in recent results (cf.[18, 19, 20, 21]). In [19], an interesting question was raised by Li as follows.
Question 1.1.
[19, Question 3.6] Let $S$ be a smooth projective surface and the closed Mori cone $\overline{NE}(S)=\sum\mathbb{R}_{\geq 0}[C_{i}]$ with $C_{i}$ some curves on $S$. Does $S$ satisfy the bounded cohomology property?
It is shown in [20, Proposition 2.6] that every smooth projective surface $S$ with $\rho(S)=1$ satisfies the bounded cohomology property. If the Picard number $\rho(S)=2$, the question 1.1 was solved for the surfaces, which satisfy either the Kodaira dimension $k(S)\leq 1$ or one of $C_{i}^{2}<0$ (cf. [18, 19, 20, 21]). Note that the $C_{i}^{2}\leq 0$ for Mori cone $\overline{NE}(S)=\sum\mathbb{R}_{\geq 0}[C_{i}]$ [21, Proposition 5.11]. Regarding the above results, we are curious to ask whether the bounded cohomology property holds for $C_{i}^{2}=0$ and $\rho(S)=2$. Thanks to the criterion for divisors, we know that if Mori cone of the fake quadrics satisfies $\overline{NE}(S)=\sum\mathbb{R}_{\geq 0}[C_{i}]$ with $C_{i}$ some curves on $S$, then $C_{i}^{2}=0$. Motivated by the above question, we consider the fake quadrics and have the following Theorem.
Theorem 1.2.
If $S$ is a fake quadric of even type, then the bounded cohomology property holds. If Mori dream space $S$ is a fake quadric of odd type, then the bounded cohomology property is true.
Note that the quadric $\mathbb{P}^{1}\times\mathbb{P}^{1}$ is a hypersurface in $\mathbb{P}^{3}$. Furthermore surface $\mathbb{P}^{2}\sharp\overline{\mathbb{P}^{2}}$ is a subvariety of $\mathbb{P}^{4}$. From this point of view, it is a natural way to ask that is there a fake quadric that can be smoothly embedded in $\mathbb{P}^{4}$? By applying the criterion for divisors (Theorem 2.2 and Theorem 3.7), we can derive the following result.
Theorem 1.3.
There is no fake quadric that can be holomorphically embedded in $\mathbb{P}^{4}$.
‘To determine whether a minimal surface with geometric genus $p_{g}=0$ is a Mori dream space’ is a very difficult problem (The definition of Mori dream space see [11, 16]). Recently, Frapporti, Keum and Lee have made some outstanding contributions to this work (cf.[11, 16]). Especially, they give a positive answer to above question for a fake quadric that is isogenous to a product of two curves (cf.[11]). For the other fake quadrics, the question is still unknown.
Hirzebruch asked whether there are fake quadrics with a trivial fundamental group (cf. [6, p.8][15, p. 780])? By the Freedman’s classification Theorem (c.f.[12]), a simply connected fake quadric, if it has existed, is homeomorphic to either $\mathbb{P}^{1}\times\mathbb{P}^{1}$ or $\mathbb{P}^{2}\sharp\overline{\mathbb{P}^{2}}$. Until now the existence of simply connected fake quadric is still unknown. All known fake quadrics are all uniformized by the bidisk, and have infinite fundamental group (cf.[27, 5, 10, 8, 9, 23] etc). This is similar to fake projective planes for which the universal cover is the unit ball in $\mathbb{C}^{2}$ (c.f. [30]).
As an application of the criterion for divisors, we give some properties of divisors with small arithmetic genus on simply connected surfaces.
Theorem 1.4.
Let $S$ be a simply-connected fake quadric. Then
  1. (i)
    $S$ is not a Mori dream space;
  2. (ii)
    the geometric genus $p_{g}(C)>2$ for any curve $C\subset S$;
  3. (iii)
    For even type (odd type), the number of curves with arithmetic genus $p_{a}\leq 5$ is at most two (three).
In [16, Question 3.2], the authors proposed whether all fake quadrics are Mori dream spaces. The Theorem 1.4 gives a negative answer to this question if the simply-connected fake quadric exists. In conclusion, we believe that this result may provide some ideas for the problem of the existence of simply-connected fake quadrics.
Acknowledgement.
We would like to thank Yifei Chen, Cheng-Yong Du, and Song Yang for their helpful communications. We would also like to extend our gratitude to Fabrizio Catanese via email for discussing the Néron–Severi lattice of the fake quadrics.
Notation and conventions
We work throughout over the complex number field $\mathbb{C}$. By a surface we mean a smooth complex projective surface, and if not specifically stated, a curve on a surface always means any non-zero effective divisor on the surface. By a negative curve we mean a reduced and irreducible curve on surface with negative self-intersection. Moreover, we use the following standard notations:
•   $h^{i}(S,\mathcal{O}_{S}(D))$: $=\dim H^{i}(S,\mathcal{O}_{S}(D))$;
•   $C.D$: the intersection number of two divisors $C$ and $D$;
•   $\iota_{\rvert D\rvert}$: a map $\iota_{\rvert D\rvert}:S\to\mathbb{P}^{\dim\rvert D\rvert}$ with $\mathcal{O}(D)=\iota_{\rvert D\rvert}^{\ast}\mathcal{O}_{\mathbb{P}^{\dim\rvert D\rvert}}(1)$;
•  $C\sim D:$ the two divisors $C,D$ are linear equivalence;
•  $C\equiv D:$ the two divisors $C,D$ are numerical equivalence;
•   $\mathrm{Nef}(S)$: the nef cone of surface $S$;
•   $\mathrm{SAmp}(S)$: the semiample cone of surface $S$.
It is noteworthy that for a simply-connected fake quadric $S$, since the torsion subgroup of $\mathrm{Pic}(S)$ is zero, then we have that $C\sim D\Leftrightarrow C\equiv D$.
2. Case I. Fake quadrics of even type
Let $S$ be a fake quadric of even type. By the definition, its Néron–Severi lattice
$\mathrm{N}^{1}(S)=\left(\begin{array}[]{cc}0&1\\ 1&0\\ \end{array}\right)$
and hence there exist two divisors $H$ and $F$ such that
$H^{2}=0,F^{2}=0\textrm{ and }H.F=1.$
From now on, we shall fix two generators $H$ and $F$ in the Néron–Severi lattice $\mathrm{N}^{1}(S)$ of a fake quadric $S$ of even type. Without loss of generality, since $K_{S}^{2}=8$, we may suppose $K_{S}\equiv 2H+2F$. For a divisor $D\equiv xH+yF$, by the Riemann–Roch theorem, we have
$\chi(\mathcal{O}_{S}(D))=(x-1)(y-1).$ (2.1)
2.1. Criterion of divisors on fake quadrics of even type: effectiveness and ampleness
Lemma 2.1.
Let $S$ be a fake quadric of even type and $C$ a reduced and irreducible curve on $S$. If $C\equiv xH+yF$ for some $x,y\in\mathbb{Z}$, then either $x\geq 0$ and $y>0$, or $x>0$ and $y\geq 0$.
Proof.
Since $K_{S}$ is ample, so $K_{S}.C>0$ and thus $x+y>0$. Using the arithmetic genus formula, we have
$p_{a}(C)=\frac{1}{2}(K_{S}.C+C^{2})+1=(x+1)(y+1)\geq 0.$ (2.2)
As a consequence, there hold only two possible cases:
  1. (1)
    $x=-1$ and $y\geq 2$, or $x\geq 2$ and $y=-1$, or
  2. (2)
    $x\geq 0$ and $y>0$, or $x>0$ and $y\geq 0$.
Next we shall rule out the case $(1)$. Suppose $C\equiv-H+yF$ with $y\geq 2$ (or $C\equiv xH-F$ with $x\geq 2$). Then its arithmetic genus $p_{a}(C)=0$ and thus $C$ is a smooth rational curve. Since $K_{S}$ is ample, then $mK_{S}$ is effective for $m\gg 0$ and hence $m(K_{S}+C)$ is effective for $m\gg 0$. By [22, Theorem 1 (2)], we have
$K_{S}.C\leq\Delta+2g(C)-2-\chi(C)=4+0-2-2=0,$
where $\Delta:=3c_{2}-c_{1}^{2}$ and $\chi(C)$ is the topological Euler characteristic of $C$. This contradicts to the ampleness of $K_{S}$ and thus finishes the proof of Lemma 2.1. ∎
Based on the above theorem, we obtain the following result.
Theorem 2.2.
Let $S$ be a fake quadric of even type and $D$ a divisor on $S$. Suppose $D\equiv xH+yF$ for some $x,y\in\mathbb{Z}$.
  1. (i)
    If $D$ is an effective divisor, then $x\geq 0$ and $y\geq 0$.
  2. (ii)
    The divisor $D$ is ample if and only if $x>0$ and $y>0$.
Proof.
(i) This is a direct consequence of Lemma 2.1. (ii) Suppose $D$ is ample. Then $D^{2}=2xy>0$ and $mD$ is effective for $m\gg 0$. As a result, it follows from (i) that $x>0$ and $y>0$. Conversely, assume $x>0$ and $y>0$ and hence $D^{2}=2xy>0$. For any irreducible curve $C\equiv aH+bF$, thanks to Lemma 2.1, there holds either $a>0,b\geq 0$ or $a\geq 0,b>0$, and thus $D.C=ay+bx>0$. Consequently, the Nakai–Moishezon criterion implies that $D$ is an ample divisor. ∎
From the Theorem 2.2 (i) and the formula (2.2), we can immediately obtain that
Corollary 2.3.
Let $S$ be a fake quadric of even type. Then for any curve $C$ on $S$, the arithmetic genus $p_{a}(C)\geq 2$.
As a consequence of the theorem above, we obtain the following result.
Corollary 2.4.
Let $S$ be a fake quadric of even type and $D$ a divisor on $S$. Suppose $D\equiv xH+yF$ for some $x,y\in\mathbb{Z}$.
  1. (i)
    If $x,y\geq 2$, then $h^{0}(S,\mathcal{O}_{S}(D))>0,$ unless $D\sim K_{S}$.
  2. (ii)
    If $x,y\geq 3$, then $h^{1}(S,\mathcal{O}_{S}(D))=h^{2}(S,\mathcal{O}_{S}(D))=0$ and
    $h^{0}(S,\mathcal{O}_{S}(D))=(x-1)(y-1)\geq 4.$
Proof.
(i) Since $K_{S}-D\equiv(2-x)H+(2-y)F$, the assumption and Theorem 2.2 (i) imply that the divisor $K_{S}-D$ is not effective. The Serre duality yields $h^{2}(S,\mathcal{O}_{S}(D))=h^{0}(S,\mathcal{O}_{S}(K_{S}-D))=0$. As a result, the Riemann–Roch theorem concludes
$h^{0}(S,\mathcal{O}_{S}(D))\geq\chi(\mathcal{O}_{S}(D))=(x-1)(y-1)\geq 1.$
(ii) By Theorem 2.2 (ii), the $x,y\geq 3$ guarantees that the divisor $D-K_{S}$ is ample. According to the Kodaira vanishing theorem, we have
$h^{1}(S,\mathcal{O}_{S}(D))=h^{2}(S,\mathcal{O}_{S}(D))=0$
and
$h^{0}(S,\mathcal{O}_{S}(D))=\chi(S,\mathcal{O}_{S}(D))=(x-1)(y-1)\geq 4.$
This completes the proof. ∎
2.2. The bounded cohomology property for fake quadrics of even type
In this subsection, a curve means a reduced and irreducible curve on a surface. Before the main result of this subsection, the following Lemmas are necessary.
Lemma 2.5.
Let $S$ be a fake quadric and $C\equiv xH+yF,x,y\in\mathbb{Z}$ a curve on $S$. If $S$ is of even type, then $C$ satisfies the bounded cohomology property for $x,y\geq 2$.
Proof.
Recall that the divisor $K_{S}-C$ is not effective. The Serre duality yields
$h^{2}(S,\mathcal{O}_{S}(C))=h^{0}(S,\mathcal{O}_{S}(K_{S}-C))=0.$
As a result, the Riemann–Roch theorem concludes
$h^{0}(S,\mathcal{O}_{S}(C))-h^{1}(S,\mathcal{O}_{S}(C))=\chi(\mathcal{O}_{S}(C))=(x-1)(y-1)>0\quad\text{for}\quad x,y\geq 2.$
Then we have that $h^{1}(S,\mathcal{O}_{S}(C))<h^{0}(S,\mathcal{O}_{S}(C)).$
Lemma 2.6.
Let $S$ be a fake quadric and $C\equiv xH+yF,x,y\in\mathbb{Z}$ a curve on $S$. If $S$ is of even type, then $C$ satisfies the bounded cohomology property for either $x=1$ or $y=1$.
Proof.
Using the same argument as above, we have that
$h^{0}(S,\mathcal{O}_{S}(C))-h^{1}(S,\mathcal{O}_{S}(C))=\chi(\mathcal{O}_{S}(C))=(x-1)(y-1)=0.$
This implies that $h^{1}(S,\mathcal{O}_{S}(C))=h^{0}(S,\mathcal{O}_{S}(C)).$ Then we have that
$h^{1}(S,\mathcal{O}_{S}(C))<2h^{0}(S,\mathcal{O}_{S}(C)).$
Proposition 2.7.
[1, Proposition 1.3] Let $S$ be a smooth projective surface with irregularity $q(x)=0$ and let $D$ be a non-trivial effective divisor of $S$ such that the linear system $\rvert D\rvert$ does not contain fixed components and $D^{2}=0$. Then $D=aC,a\in\mathbb{Z}_{>0}$ with $C$ smooth and integral, $h^{0}(S,\mathcal{O}_{S}(C))=2$ and $H^{0}(S,\mathcal{O}_{S}(D))\cong\mathrm{Sym}^{a}H^{0}(S,\mathcal{O}_{S}(C))$.
Lemma 2.8.
Let $S$ be a fake quadric and $C\equiv xH+yF,x,y\in\mathbb{Z}$ a curve on $S$. If $S$ is of even type, then $C$ satisfies the bounded cohomology property for either $x=0$ or $y=0$.
Proof.
By Proposition 2.7, we have that $h^{0}(S,\mathcal{O}_{S}(C))\leq 2$ and there exists a positive integral $k_{0}$ such that the curve $C\sim kH$ or $kF$ for $k\leq k_{0}$. According to
$h^{0}(S,\mathcal{O}_{S}(C))-h^{1}(S,\mathcal{O}_{S}(C))=\chi(\mathcal{O}_{S}(C))=(x-1)(y-1)=-(k-1),$
we have that
$h^{1}(S,\mathcal{O}_{S}(C))/h^{0}(S,\mathcal{O}_{S}(C))=[h^{0}(S,\mathcal{O}_{S}(C))+(k-1)]/h^{0}(S,\mathcal{O}_{S}(C))$
$\leq[h^{0}(S,\mathcal{O}_{S}(C))+(k_{0}-1)]/h^{0}(S,\mathcal{O}_{S}(C)).$
Define that $c:=[h^{0}(S,\mathcal{O}_{S}(C))+(k_{0}-1)]/h^{0}(S,\mathcal{O}_{S}(C))+1$. Then we have that
$h^{1}(S,\mathcal{O}_{S}(C))<ch^{0}(S,\mathcal{O}_{S}(C)).$
Now we begin to prove the Theorem 1.2
Proof of Theorem 1.2 of even type..
Let $c_{S}:=max\{2,c\}$. Then we have that
$h^{1}(S,\mathcal{O}_{S}(C))<c_{S}h^{0}(S,\mathcal{O}_{S}(C)).$
2.3. Embedding fake quadrics of even type in $\mathbb{P}^{4}$
It is well-known that every smooth projective surface is holomorphically embedded in $\mathbb{P}^{5}$. However, a smooth projective surface may not in general be embedded in $\mathbb{P}^{4}$. As a by-product, we have the following observation which we believe is new.
Proof of Theorem 1.3 of even type..
Let $S$ be a fake quadric and $\iota:S\hookrightarrow\mathbb{P}^{4}$ an embedding. Then there exists a very ample divisor $D$ such that $\mathcal{O}_{S}(D)=\iota^{\ast}\mathcal{O}_{\mathbb{P}^{4}}(1)$. Moreover, the fake quadric $S$ satisfies the following numerical condition (c.f. [14, Appendix A, Example 4.1.3]):
$d^{2}-10d-5D.K_{S}-2K_{S}^{2}+12+12p_{a}(S)=0,$ (2.3)
where $d=D^{2}$ and the arithmetic genus $p_{a}(S)=0$. Assume that $D\equiv xH+yF$. It follows from Theorem 2.2 that $x>0$ and $y>0$. Therefore, the equation (2.3) implies that $x$ and $y$ must satisfy
$02x^{2}y^{2}-10xy-5x-5y-2=0,\quad\text{for}\quad x>0,y>$ 0 (2.4)
By [33, Chapter V, Corollary 2.9], we have that $h^{0}(S,\mathcal{O}_{S}(D))<\binom{4}{2}=6$. Together with Corollary 2.4 (ii), we have that
$(x-1)(y-1)=h^{0}(S,\mathcal{O}_{S}(D))<6,\quad\text{for}\quad x\geq 3,y\geq 3.$ (2.5)
If $30<x<$ 3 or $30<y<$ 3, there is no solution of equation (2.4). For $x\geq 3,y\geq 3$, combine with (2.4) and (2.5), it can be concluded that the values of $x,y$ do not exist. This completes the proof. ∎
2.4. Curves on simply-connected fake quadrics of even type
In this section, we discuss some properties of curves on simply-connected fake quadrics. Before the proof of main result, some useful Lemmas are needed.
Lemma 2.9.
[4] Let $f:S\to\mathbb{P}^{2}$ be a double cover between two algebraic surfaces $S$ and $\mathbb{P}^{2}$. If $K_{S}^{2}=8$ and $p_{g}(S)=0$, then the Kodaira dimension of $S$ is $kod(S)=-\infty$.
Proof.
Suppose that the double cover $f:S\to\mathbb{P}^{2}$ is branched along divisor $B$ and $B=2L$ for some $L\in\mathrm{Pic}(\mathbb{P}^{2})$. Let $L=mH$, where $H$ is a hyperplane of $\mathbb{P}^{2}$. Recall from [4, V.22] that
$p_{g}(S)=1+\frac{1}{2}m(m-3),\quad K_{S}^{2}=2(m-3)^{2}.$
So by the assumption that $m=1.$ Since
$K_{S}=f^{\ast}(K_{\mathbb{P}^{2}}+L)=f^{\ast}(K_{\mathbb{P}^{2}}+H)=-2f^{\ast}(H)$
then the canonical divisor $K_{S}$ is anti-ample. Indeed, since $f:S\to\mathbb{P}^{2}$ is a finite morphism, by [17, Corollary 1.2.15], then $f^{\ast}(H)$ is ample. This ensures that $K_{S}$ is anti-ample. Thus the result holds. ∎
Lemma 2.10.
If $\pi:S\to C$ is a fibration with irregularity $q(S)=0$, then the genus $g(C)=0$ and $C\cong\mathbb{P}^{1}$.
Proof.
Assume that $g(C)\neq 0$. Let $\alpha\in H^{1,0}(C)$ be a nonzero element. Then we have that $\pi^{\ast}\alpha\neq 0$. This shows that $H^{1,0}(S)\neq 0$. This contradicts the irregularity $q(S)=0$. ∎
Lemma 2.11.
Let $S$ be a fake quadric and $D\subset S$ be a divisor with $\dim\rvert D\rvert>0$. If all elements in the linear system $\rvert D\rvert$ are reduced and irreducible, then $K_{S}+D$ is very ample when the arithmetic genus $p_{a}(D)\geq 3$.
Proof.
Suppose that $D$ is effective. Let $\iota:=\iota_{\rvert K_{S}+D\rvert}.$ By the exact sequence
$0\to\mathcal{O}_{S}(K_{S})\to\mathcal{O}_{S}(K_{S}+D)\to\mathcal{O}_{D}(K_{D})\to 0$
we see that $h^{0}(S,\mathcal{O}_{S}(K_{S}+D))\to h^{0}(D,\mathcal{O}_{D}(K_{D}))$ is surjective. This ensures that the canonical map $\iota_{K_{D}}:D\to\mathbb{P}^{g-1}$ is essentially the restriction map $\iota\rvert_{D}$, where $g:=p_{a}(D)$ is the arithmetic genus of $D$.
Note that the degree of the canonical map $\iota\rvert_{D}:D\to\iota(D)\subset\mathbb{P}^{g-1}$ is either $1$ or $2$ for $p_{a}(D)\geq 3$, and if $\deg(\iota\rvert_{D})=1$, then it is an embedding, if $\deg(\iota\rvert_{D})=2$, then $\iota(D)=\mathbb{P}^{1}$. Since $\dim\rvert D\rvert>0$, then $S$ is covered by any pencil $L\subset\rvert D\rvert$. This implies that if $\deg\iota\rvert_{D}=1$, then $\iota:S\to\mathbb{P}^{\dim\rvert K_{S}+D\rvert}$ is an embedding; if $\deg\iota\rvert_{D}=2$, then $\deg\iota=2$. Indeed, the $\deg(\iota\rvert_{D})=1,2$ guarantees that $\deg(\iota\rvert_{D^{\prime}})=1,2$ for $D,D^{\prime}\in L$. Otherwise, assume that $\deg\iota\rvert_{D}=1$ and $\deg\iota\rvert_{D^{\prime}}=2$. Blowing up the base points of $L$ step by step, together with Lemma 2.10, we get a fibration $B$ with base space $\mathbb{P}^{1}$ and fibers that are all elements of $L$ (To be precise, all fibers of $B$ correspond one-to-one with all elements of $L$. To make it easier to explain, we regard them as the same one). Then there is a fibre mapping $\bar{\iota}$ on $B$, which induced by $\iota$, such that the fibers of fibration $\bar{\iota}(B)$ contain $D$ and $2\mathbb{P}^{1}$ (The reason is that $\bar{\iota}_{\ast}:D\to D$ and $\bar{\iota}_{\ast}:D^{\prime}\to 2\mathbb{P}^{1}$). This gives that
$p_{a}(D)=p_{a}(2\mathbb{P}^{1})=2p_{a}(\mathbb{P}^{1})+\mathbb{P}^{1}.\mathbb{P}^{1}-1=\mathbb{P}^{1}.\mathbb{P}^{1}-1=-1$
a contradiction. Therefore, we have that
$\deg\iota=\deg\bar{\iota}=\deg\bar{\iota}_{D}=\deg\iota_{D}$
for any $D\in L$ (more details about $\deg\bar{\iota}=\deg\bar{\iota}_{D}$ see following Remark).
If $\deg\iota=2$, by $\bar{\iota}:D\to\mathbb{P}^{1}$, then $\bar{\iota}(B)$ is a fiberation with fibers $\mathbb{P}^{1}$ and base space $\mathbb{P}^{1}$. This ensures that $\bar{\iota}(B)$ is either $\mathbb{P}^{1}\times\mathbb{P}^{1}$ or $\mathbb{P}^{2}\sharp\overline{\mathbb{P}^{2}}$ (the Hirzebruch surfaces). From the following diagram
the $\bar{\iota}(B)$ can be constructed by blowing up the base points of the pencil $\iota_{\ast}L$ step by step. This means that $\bar{\iota}(B)=\mathbb{P}^{2}\sharp\overline{\mathbb{P}^{2}}$. By using the above diagram again, we have that $\iota(S)=\pi(\mathbb{P}^{2}\sharp\overline{\mathbb{P}^{2}})=\mathbb{P}^{2}.$ This contradicts Lemma 2.9.
Therefore, the restriction map $\iota\rvert_{D}:D\to\mathbb{P}^{g-1}$ is an embedding. Then we have that $\iota:S\to\mathbb{P}^{\dim\rvert K_{S}+D\rvert}$ also is an embedding. This gives that the divisor $K_{S}+D$ is very ample. ∎
Remark 2.12.
There is an important detail to be noted for the proof of $\deg\bar{\iota}\rvert_{F}=\deg\bar{\iota}$ in Lemma 2.11. To be specific, we need to demonstrate that the restriction of $\bar{\iota}$ on base spaces, $\bar{\iota}:\mathbb{P}^{1}\to\mathbb{P}^{1}$, is an isomorphism. Thus we will explain that the degree of the map $\bar{\iota}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ is $1$. This allows us to show that the map $\bar{\iota}$, which is restricted to any section of $B$, has degree $1$. Suppose that $B=S\sharp_{k}\overline{\mathbb{P}^{2}}$ and $\bar{\iota}(B)=\iota(S)\sharp_{l}\overline{\mathbb{P}^{2}}$. By the following diagram
${S\sharp_{k}\overline{\mathbb{P}^{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\bar{\iota}}$$\scriptstyle{\pi}$${\iota(S)\sharp_{l}\overline{\mathbb{P}^{2}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\pi}$${S\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\iota}$${\iota(S)}$
we have that the restriction map
$\text{from}\quad\text{any}\quad\overline{\mathbb{P}^{2}}-D^{2}\subset S\sharp_{k}\overline{\mathbb{P}^{2}}\quad\text{ to}\quad\overline{\mathbb{P}^{2}}-D^{2}\subset\iota(S)\sharp_{l}\overline{\mathbb{P}^{2}}$
has degree $1$. Note that any section of $B$ is contained in
$\text{any}\quad\overline{\mathbb{P}^{2}}-D^{2}\subset S\sharp_{k}\overline{\mathbb{P}^{2}},i.e.\mathbb{P}^{1}\subset\overline{\mathbb{P}^{2}}-D^{2}.$
Thus the degree of $\bar{\iota}:\mathbb{P}^{1}\to\mathbb{P}^{1}$ is $1$.
With the above preparation, we now turn to the central focus of this subsection.
Lemma 2.13.
Let $S$ be a fake quadric of even type, and $C\subset S$ a reduce and irreducible curve. If $C\sim mH$ or $C\sim mF$ for some $m\in\mathbb{Z}_{>0}$, then $C$ is smooth.
Proof.
Without loss of generality, we assume that $C\sim mH$. By
$2m=(2H+2F).mH=K_{S}.C\leq\Delta+2g(C)-2-\chi(C)\leq\Delta+2p_{a}(C)-2-\chi(C)=2(m+1)-2=2m$
we have that the arithmetic genus $p_{a}(C)=p_{g}(S)=m+1.$ This gives that $C$ is smooth. ∎
This result implies that if $S$ is a fibration $f:S\to\mathbb{P}^{1}$, every singular fiber is multiple fibers with smooth support.
Lemma 2.14.
Let $S$ be a simply-connected fake quadric of even type. If a divisor $C\subset S$ satisfies $C\sim H$ or $C\sim F$ or $C\sim H+F$, then $h^{0}(S,\mathcal{O}_{S}(C))=0$.
Proof.
Since the geometric genus $p_{g}(S)=0$, then the canonical divisor $K_{S}$ is not effective. If $C\sim H$ is an effective divisor, by $K_{S}\sim 2H+2F$, we have that the divisors, which are linearly equivalent to $H+2F$ or $H+F$ or $F$, are not effective. Under the above discussion, we have that any divisor $D\in\rvert 3H+2F\rvert$ cannot be decomposed into the sum of two other divisors and so $D$ is reduced and irreducible.
According to Lemma 2.4 (i), we conclude that $\dim\rvert D\rvert>0.$ From the formula (2.2), the arithmetic genus $p_{a}(D)=(3+1)(2+1)=12.$ Let $E:=K_{S}+D$. The Lemma 2.11 gives that $E$ is very ample on $S$. Thus $E\rvert_{C}$ is very ample on $C$. Thanks to $\deg_{C}E=C\cdot E=4>2g(C)-2=2$, we have that
$h^{0}(C,\mathcal{O}_{C}(E))=\chi(C,\mathcal{O}_{C}(E))=\deg_{C}E+1-p_{a}(C)=3.$
This implies that $C$ is a plane curve with degree $d=C\cdot E=4$. Thus we conclude that the arithmetic genus $p_{a}(C)=\frac{(d-1)(d-2)}{2}=3$. This contradicts that $p_{\alpha}(C)=2$.
Applying the same argument as that used in the proof of the case $C\sim H$, we can demonstrate that the assertion also holds for $C\sim F$ by $D\in\rvert 2H+3F\rvert$. ∎
Proof of Theorem 1.4 of even type..
(i) Note that if $S$ is a Mori dream space, then $S$ is a fibration $f:S\to\mathbb{P}^{1}$. Thanks to Theorem 2.2, the fiber of such fibration is $\mathbb{F}\sim mH$ or $\mathbb{F}\sim mF$ for some $m\mathbb{Z}_{>0}.$ It is shown in [29] that if $S$ is simply-connected, then the singular fibers of the fibration are not multiple fibers. Together with Lemma 2.13, we give that $S$ is a fiber bundle over $\mathbb{P}^{1}.$ From the homotopy exact sequence
$\ldots\pi_{2}(\mathbb{F})\to\pi_{2}(S)\to\pi_{2}(\mathbb{P}^{1})\to\pi_{1}(\mathbb{F})\to\pi_{1}(S)\to\ldots$
we conclude that the fundamental group $\pi_{1}(S)$ is nontrivial, a contradiction.
(ii) Thanks to the inequality $K_{S}.C\leq\Delta+2g(C)-2-\chi(C)$ again, together with $K_{S}.C>0$, we have that the geometric genus $p_{g}(C)\geq 2.$ If $p_{g}(C)=2$, then $K_{S}.C\leq 2.$ From Theorem 2.2 (i), the
$(2H+2F).(xH+yF)=K_{S}.C\leq 2$
shows that either $x=1,y=0$ or $x=0,y=1.$ Therefore, the curve $C\sim H$ or $C\sim F$. The Lemma 2.13 provides that $g_{a}(C)=p_{a}(C)=2$. This contradicts Lemma 2.14.
(iii) According to Theorem 2.2 (i), we only consider $C\sim xH+yF$ with $x,y\geq 0$. By the arithmetic genus $p_{\alpha}(C)=(x+1)(y+1)$, together with Lemma 2.14, we conclude that
  1. (a)
    if $p_{a}(C)=3$, then $C\sim 2H$ or $C\sim 2F$;
  2. (b)
    if $p_{a}(C)=4$, then $C\sim 3H$ or $C\sim 3F$;
  3. (c)
    if $p_{a}(C)=5$, then $C\sim 4H$ or $C\sim 4F$;
Recall that $S$ is not a fibration. This gives that
$h^{0}(S,\mathcal{O}_{S}(mD))\leq 1\quad\text{for}\quad m\in\mathbb{Z}_{>0},\quad D\sim H\quad\text{ or}\quad D\sim F.$
This finishes the proof. ∎
3. Case II. Fake quadrics of odd type
Let $S$ be a fake quadric of odd type. According to the definition, the Néron–Severi lattice
$\mathrm{N}^{1}(S)=\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$
and thus there are two divisors $H$ and $F$ with intersection numbers
$H^{2}=1,F^{2}=-1\textrm{ and }H.F=0.$
From now on, we will fix two generators $H$ and $F$ of the Néron–Severi lattice $\mathrm{N}^{1}(S)$. Note that the canonical divisor $K_{S}$ is a primitive divisor in $\mathrm{N}^{1}(S)$ (c.f. [28, Theorems 3.1 (iii), 3.10]). Without loss of generality, since $K_{S}^{2}=8$, we may assume $K_{S}\equiv 3H-F$. For a divisor $D\equiv xH+yF$, by the Riemann–Roch theorem, we have
$\chi(\mathcal{O}_{S}(D))=\frac{1}{2}(x+y-1)(x-y-2).$ (3.1)
3.1. Criterion of divisors on fake quadrics of odd type: effectiveness and ampleness
Lemma 3.1.
Let $S$ be a fake quadric of odd type and $C$ a reduced and irreducible curve on $S$. If $C\equiv xH+yF$ for some $x,y\in\mathbb{Z}$, then $x\geq 0$ and $-x-1<y\leq x+1$.
Proof.
Since $K_{S}$ is ample, then $K_{S}.C>0$ and thus $03x+y>$ 0. By the arithmetic genus formula, we get
$p_{a}(C)=\frac{1}{2}(x+y+1)(x-y+2)\geq 0.$ (3.2)
Then we have only two possible cases:
  1. (1)
    $x=0$, $y=2$; $x>0$ and $y=-(x+1)$; $x>0$ and $y=x+2$, or
  2. (2)
    $x\geq 0$ and $-(x+1)<y\leq x+1$.
We will rule out the case $(1)$. The idea is as same as the proof of Lemma 2.1. Suppose $C\equiv 2F$ or $C\equiv xH-(x+1)F$ or $C\equiv xH+(x+2)F$ with $x>0$. Then the arithmetic genus $p_{a}(C)=0$ and thus $C$ is a smooth rational curve. Since $K_{S}$ is ample, so $mK_{S}$ is effective for $m\gg 0$ and thus $m(K_{S}+C)$ is effective for some $m\gg 0$. By [22, Theorem 1 (2)], we have $K_{S}.C\leq\Delta+2g(C)-2-\chi(C)=4+0-2-2=0$. This contradicts to the ampleness of $K_{S}$. ∎
To more accurately characterize the divisors, in the following we mainly focus on the existence of negative curves on fake quadrics.
Lemma 3.2.
If the Néron–Severi lattice of fake quadric $S$ is odd type, then there is at most one negative curve on $S$.
Proof.
According to Lemma 3.1, any negative curve (reduced and irreducible) is numerically equivalent to $xH+(x+1)F$ for some $x\in\mathbb{Z}_{\geq 0}$. Suppose that $x_{0}\in\mathbb{Z}_{\geq 0}$ is the minimal number such that $x_{0}H+(x_{0}+1)F$ is numerically equivalent to an effective divisor. Let curve $C\equiv x_{0}H+(x_{0}+1)F$. The $C^{2}<0$ gives that the linear system $\rvert C\rvert$ contains a fixed component. This implies that $\dim\rvert C\rvert=0.$
For any divisor $D\equiv xH+(x+1)F,x>x_{0}$, we may conclude that if $D$ effective, then $D$ is not reduced or irreducible. Otherwise, then $\dim\rvert D\rvert=0$ and so $D-C$ is not effective. By $\deg D\rvert_{C}<0,$ we have $h^{0}(C,\mathcal{O}_{C}(D))=0.$ According to the exact sequence
$0\to h^{0}(S,\mathcal{O}_{S}(D-C))\to h^{0}(S,\mathcal{O}_{S}(D))\to h^{0}(C,\mathcal{O}_{C}(D))\to\ldots$
together with $h^{0}(S,\mathcal{O}_{S}(D-C))=0,$ we have $h^{0}(S,\mathcal{O}_{S}(D))=0$, a contradiction. Thus any negative curve is numerically equivalent to $C$.
Suppose that $C^{\prime}\equiv C$ is another negative curve. Then $C$ and $C^{\prime}$ only differ by a torsion divisor. This shows that there exists a $m\in\mathbb{Z}_{>0}$ such that $mC\sim mC^{\prime}$. Thus we that $h^{0}(S,\mathscr{O}_{S}(mC))\geq 2$. Since $(kC)^{2}<0$ for any $k\leq m$, then $h^{0}(S,\mathscr{O}_{S}(mC))=1$, a contradiction. ∎
As an application, we have that
Corollary 3.3.
Suppose that the Néron–Severi lattice of fake quadric $S$ is the odd type. Let $C\equiv x_{0}H+(x_{0}+1)F$ be the unique negative curve on $S$. If a divisor $D$ on $S$ is nef and big but not ample, then
[…]
Proof.
As stated in Lemma 3.1, together with Lemma 3.2, the boundary of effective cone of $S$ is generated by $C$ and $H-F.$ If $D$ is nef but not ample, then
$\text{either}\quad D.C=0\quad\text{or}\quad D.(H-F)=0.$
This shows that
$\text{either}\quad D\equiv m(H-F)\quad\text{ or}\quad D\equiv m[(x_{0}+1)H+x_{0}F]\quad\text{ for}\quad m\geq 0.$
Since $D$ is big, $D^{2}\neq 0$. Thus the result holds. ∎
Definition 3.4.
[25] A projective variety $X$ is said to be $\mathbb{Q}$-Gorenstein, if $X$ is Cohen-Macaulay and the canonical divisor $K_{X}$ is a $\mathbb{Q}$-Cartier divisor.
The following result is very helpful for the proof of the Lemma 3.6.
Proposition 3.5.
[4, III. (2.1) Theorem] [Grauert’s criterion] A reduced, compact connected curve $C$ with irreducible components $C_{i}$ on a smooth surface is exceptional if and only if the intersection matrix $(C_{i}C_{j})$ is negative definite. Furthermore, a surface obtained by contracting curve $C$ to a point is a normal surface.
With the above preparations, we begin to consider Lemma 3.6. The following result is the core part of this subsection.
Lemma 3.6.
If the Néron–Severi lattice of fake quadric $S$ is the odd type, the negative curve does not existed.
Proof.
If negative curve $C\subset S$ exists, the Lemma 3.2 implies that it is unique in the numerical equivalence class of $x_{0}H+(x_{0}+1)F$ for $x_{0}\in\mathbb{Z}_{\geq 0}$. Consider the morphism $f:S\to S^{\prime}$, which only contracts $C$ to a point. Since $C^{2}<0$, according to Grauert’s criterion, we may conclude that $S^{\prime}$ is a normal surface. By the fact that every normal surface is Cohen-Macaulay, then we have that $S^{\prime}$ is Cohen-Macaulay.
In the following, we need show that $S^{\prime}$ is a $\mathbb{Q}$-Gorenstein variety. This makes us to prove that the canonical divisor $K_{S^{\prime}}$ of $S^{\prime}$ is a $\mathbb{Q}$-Cartier divisor. Denote that $D:=f_{\ast}H$. By $f_{\ast}[x_{0}H+(x_{0}+1)F]\equiv f_{\ast}C=0,$ we have that $f_{\ast}F\equiv-\frac{x_{0}}{x_{0}+1}D$. Note that the canonical sheaf of $S^{\prime}$ is defined by $\omega_{S^{\prime}}=i_{\ast}\omega_{S^{\prime}_{0}}$, where $S_{0}^{\prime}$ is the regular part of $S^{\prime}$ and $i:S^{\prime}_{0}\to S^{\prime}$ is an inclusion map. From the Levi’s extension Theorem [4, I.(8.7) Theorem] or Riemann’s extension Theorem [13], this implies that the canonical divisor
$K_{S^{\prime}}=f_{\ast}K_{S}=3f_{\ast}H-f_{\ast}F\equiv 3D-\frac{x_{0}}{x_{0}+1}D=\frac{2x_{0}+3}{x_{0}+1}D.$
Let $L$ be a divisor such that $f=\iota_{\rvert L\rvert}$. Recall that $L$ is ample if and only if the morphism $\iota_{\rvert L\rvert}$ is a finite morphism [17, Corollary 1.2.15]. This gives that the divisor $L$ is nef and big but not ample. According to Corollary 3.3, we have that $L\equiv m[(x_{0}+1)H+x_{0}F]$ for $m\in\mathbb{Z}_{>0}$. Since the second betti number of $S^{\prime}$ is $1$, we can assume that
$m[(x_{0}+1)H+x_{0}F]\equiv L\equiv f^{\ast}(rD),\quad r\in\mathbb{R}.$
By the fact that the generically degree of $f$ is $1$ ($f:S-C\to S^{\prime}$ has degree $1$), we have $D=f_{\ast}f^{\ast}D$. This indicates that
$D\equiv$ $\frac{m}{r}f_{\ast}[(x_{0}+1)H+x_{0}F]=m\frac{(x_{0}+1)f_{\ast}H+x_{0}f_{\ast}F}{r}$
$\equiv$ $m\frac{x_{0}+1}{r}D-m\frac{x_{0}^{2}}{r(x_{0}+1)}D=\frac{m(2x_{0}+1)}{r(x_{0}+1)}D$
and then $r=m\frac{2x_{0}+1}{x_{0}+1}\in\mathbb{Q}$. For convenience, assume that the $S^{\prime}\subset\mathbb{P}^{\dim\rvert L\rvert}$ is determined by the embed map $i:S^{\prime}\hookrightarrow\mathbb{P}^{\dim\rvert L\rvert}$. Thanks to
$m[(x_{0}+1)H+x_{0}F]\equiv f^{\ast}\mathcal{O}_{\mathbb{P}^{\dim\rvert L\rvert}}(1),$
we have that $rD\equiv i^{\ast}\mathcal{O}_{\mathbb{P}^{\dim\rvert L\rvert}}(1)$. Thus $rD$ is Cartier divisor and so $D$ is $\mathbb{Q}$-Cartier divisor. This reveals that the canonical divisor $K_{S^{\prime}}\equiv\frac{2x_{0}+3}{x_{0}+1}D$ also is a $\mathbb{Q}$-Cartier divisor. Thus $S^{\prime}$ is $\mathbb{Q}$-Gorenstein. Note that the singular point of $\mathbb{Q}$-Gorenstein variety is rational singularity (c.f.[25]). As a result, the singular point of $S^{\prime}$ is the rational singularity.
According to Artin’s criterion[2, Lemma 3.8], we conclude that the arithmetic genus $p_{a}(C)=0$. This contradicts that
$p_{a}(C)=\frac{1}{2}[x_{0}+(x_{0}+1)+1][x_{0}-(x_{0}+1)+2]=x_{0}+1>0.$
Therefore, the result holds. ∎
From the above results, it can be inferred that
Theorem 3.7.
Let $S$ be a fake quadric of odd type and $D$ a divisor on $S$. Assume $D\equiv xH+yF$ for $x,y\in\mathbb{Z}$.
  1. (i)
    If $D$ is an effective divisor, then $x\geq\rvert y\rvert$.
  2. (ii)
    If $D$ is ample if and only if $x>\rvert y\rvert$
Proof.
(i) It follows immediately from Lemma 3.1 and Lemma 3.6. (ii) Suppose $D$ is ample. Then $D^{2}=x^{2}-y^{2}>0$ and $mD$ is effective for $m\gg 0$. As a result, it follows from (i) that $x>\rvert y\rvert$. Conversely, assume $x>\rvert y\rvert$ and hence $D^{2}=x^{2}-y^{2}>0$. For any irreducible curve $C\equiv aH+bF$, thanks to (i), thus $D.C=ax-by>0$. Consequently, the Nakai–Moishezon criterion implies that $D$ is an ample divisor. ∎
Remark 3.8.
If we add $\frac{1}{2}(H+F)$ to the Néron–Severi lattice of odd type, then Néron–Severi lattice from the even type to the odd type under the following map $f$
$f:H\to H-F,\quad f:F\to\frac{1}{2}(H+F)$
is injective. This means that the properties of $H-F$ and $H+F$ in the odd type are similar to those of $H$ and $2F$ in the even type.
By the result of Beauville [7, (4.1),(4.2)], for a fake quadric $S$ of unmixed type, the tangent bundle $T_{S}$ splits as a direct sum of two line bundles $L_{1},L_{2}$ such that
$K_{S}\equiv L_{1}+L_{2},\quad L_{1}^{2}=L_{2}^{2}=0,\quad c_{1}^{2}(S)=2L_{1}.L_{2}=2c_{2}(S).$
Recall that for the odd type, the canonical divisor is $K_{S}\equiv 3H-F$. This implies that $L_{1}\equiv H+F,L_{2}\equiv 2(H-F)$. So even in the fake quadrics of unmixed type, the fake quadrics of odd type cannot be excluded by using the current approach. This further indicates that the odd type should be seriously considered in the study of fake quadrics.
As a by-product, we have the following Corollary.
Corollary 3.9.
Let $S$ be a fake quadric of odd type. Then for any curve $C$ on $S$, the arithmetic genus $p_{a}(C)\geq 2$.
As another consequence of theorem 3.7, we can obtain the following result.
Corollary 3.10.
Let $S$ be a fake quadric of odd type and $D$ a divisor on $S$. Suppose $D\equiv xH+yF$.
  1. (i)
    If $x+y>1$ and $x-y>2$, then $h^{0}(S,\mathcal{O}_{S}(D))>0$ unless $D\sim K_{S}$;
  2. (ii)
    If $x-3>\rvert y+1\rvert$, then $h^{1}(S,\mathcal{O}_{S}(D))=h^{2}(S,\mathcal{O}_{S}(D))=0$ and
    $h^{0}(S,\mathcal{O}_{S}(D))=\frac{1}{2}(x+y-1)(x-y-2)\geq 3.$
Proof.
(i) If $K_{S}-D\equiv(3-x)H-(y+1)F$, the assumptions and Theorem 3.7 imply that the divisor $K_{S}-D$ is not effective. Indeed, by $x+y>1$ and $x-y>2$, we have that $x\geq 3$ and $x\geq\rvert y\rvert$. This gives that $(3-x)<\rvert y+1\rvert$ unless $D\equiv 3H-F$. By the Serre duality and Theorem 3.7 (i), we obtain
$h^{2}(S,\mathcal{O}_{S}(D))=h^{0}(S,\mathcal{O}_{S}(K_{S}-D))=0$
As a result, the Riemann–Roch theorem implies
$h^{0}(S,\mathcal{O}_{S}(D))\geq\chi(S,\mathcal{O}_{S}(D))=\frac{1}{2}(x+y-1)(x-y-2)\geq 1.$
(ii) By assumption, the divisor $D-K_{S}=(x-3)H+(y+1)F$ satisfies $x-3>\rvert y+1\rvert$. Hence, Theorem 3.7 implies that the divisor $D-K_{S}$ is ample. Moreover, according to the Kodaira vanishing theorem, we have
$h^{i}(S,\mathcal{O}_{S}(D))=0$
for $i>0$. The Riemann–Roch theorem yields
$h^{0}(S,\mathcal{O}_{S}(D))=\chi(S,\mathcal{O}_{S}(D))=\frac{1}{2}(x+y-1)(x-y-2)\geq 3.$
This completes the proof. ∎
3.2. The bounded cohomology property for fake quadrics of odd type
In this subsection, a curve also means a reduced and irreducible curve on a surface. Using the same argument as the even type (Lemma 2.5, Lemma 2.6 and Lemma 2.8), we can also prove the following results.
Lemma 3.11.
Let $S$ be a fake quadric and $C\equiv xH+yF,x,y\in\mathbb{Z}$ a curve on $S$. If $S$ is of odd type, then $C$ satisfies the bounded cohomology property for $x+y>1$ and $x-y>2$.
Lemma 3.12.
Let $S$ be a fake quadric and $C\equiv xH+yF,x,y\in\mathbb{Z}$ a curve on $S$. If $S$ is of odd type, then $C$ satisfies the bounded cohomology property for either $x+y=1$ or $x-y=2$.
Lemma 3.13.
Let $S$ be a fake quadric and $C\equiv xH+yF,x,y\in\mathbb{Z}$ a curve on $S$. If $S$ is of odd type, then $C$ satisfies the bounded cohomology property for $x=\pm y$.
However, if $C$ satisfies $x=y+1$, then
$h^{0}(S,\mathcal{O}_{S}(C))-h^{1}(S,\mathcal{O}_{S}(C))=\chi(\mathcal{O}_{S}(C))=\frac{1}{2}(x+y-1)(x-y-2)=-y.$
This implies that the bounded cohomology property can hardly be determined for
$C\equiv(x+1)H+xF.$
Therefore we only consider that $S$ is a Mori dream space.
Before the proof of the Theorem 1.2 of odd type, a useful result is required in the following.
Lemma 3.14.
[21, Lemma 4.3] Let $S$ be a smooth projective surface and the $\overline{NE}(S)=\sum\mathbb{R}_{\geq 0}[C_{i}]$ with some curves $C_{i}$ on $S$. Suppose that the kodaira dimension of $C_{1},C_{2}$ satisfy
$k(S,C_{1})=k(S,C_{2})=1.$
Then $S$ satisfies the bounded cohomology property.
With the above preparation, we begin to prove the Theorem 1.2
Proof of Theorem 1.2 of the odd type..
If $S$ is a Mori dream space, by [11, Corollary 1.8]or [16, Corollary 2.10], we have
$\mathrm{Nef}(S)=\mathrm{SAmp}(S)=\overline{NE}(S).$
The same notations as in Lemma 3.14, this gives that the curves
$C_{1}\equiv k_{1}(H-F)\quad\text{and}\quad C_{2}\equiv k_{2}(H+F)$
for some $k_{1},k_{2}\in\mathbb{Z}_{\geq 0}$. By Proposition 2.7, we have that
$k(S,C_{1})=k(S,C_{2})=1.$
According to the Theorem 3.14, the result holds. ∎
3.3. Embedding fake quadrics of odd type in $\mathbb{P}^{4}$
In this subsection, we continue to study the fake quadrics in $\mathbb{P}^{4}$ and finish the remaining proof of the Theorem 1.3.
Proof of Theorem 1.3 of odd type..
The same notation as the even type, assume that $D\equiv xH+yF$ is an ample divisor. Therefore, the equation (2.3) implies that $x$ and $y$ must satisfy
$(x^{2}-y^{2})^{2}-10(x^{2}-y^{2})-5(3x+y)-4=0.$
It follows from Theorem 3.7 (ii) that $x>\rvert y\rvert$. By suitably modifying the previous equation, we can obtain that
$(x^{2}-y^{2}-5)^{2}=5(3x+y)+29,\quad\text{for}\quad x>\rvert y\rvert$ (3.3)
If $x+y>1$ and $x-y>2$, by Corollary 3.10 (i), we conclude that
$\frac{1}{2}(x+y-1)(x-y-2)=\chi(S,\mathcal{O}_{S}(D))\leq h^{0}(S,\mathcal{O}_{S}(D))<6$
unless $D\sim K_{S}$. This gives that
$(x+y-1)(x-y-2)\leq 10,\quad\text{for}\quad x+y>1,\quad x-y>2$ (3.4)
except for $x=3,y=-1$.
If $x+y\leq 1$ or $x-y\leq 2$ for $x>\rvert y\rvert$, there is no solution of equation (3.3). For $x+y>1$ and $x-y>2$, based on equations (3.3) and (3.4), it can be concluded that the values of $x,y$ are non-existent. Thus we finish the proof. ∎
3.4. Curves on simply-connected fake quadrics of odd type
Before we finish the remaining proof of the Theorem 1.4, the following results are required.
Lemma 3.15.
Let $S$ be a fake quadric of odd type, and $C\subset S$ a reduce and irreducible curve. If $C\sim m(H+F)$ or $C\sim m(H-F)$ for some $m\in\mathbb{Z}_{>0}$, then $C$ is smooth.
Proof.
Using the same argument as in Theorem 2.13, the result is true. ∎
The following result is the core content of this subsection.
Lemma 3.16.
Let $S$ be a simply-connected fake quadric of odd type, and a divisor $C\subset S$.
  1. (i)
    if the curve $C\sim H-F$, then $h^{0}(S,\mathcal{O}_{S}(C))=0$.
  2. (ii)
    furthermore,
[…]
Proof.
(i) Since the geometric genus $p_{g}(S)=0$, then the canonical divisor $K_{S}$ is not effective. If $C\sim H-F$ is an effective divisor, by $K_{S}\sim 3H-F$, we have that the divisors, which are linearly equivalent to $H$ or $2H$, are not effective. Under the above discussion, we have that any curve $D\in\rvert 4H-2F\rvert$ cannot be decomposed into the sum of two other divisors and so $D$ is reduced and irreducible.
According to Corollary 3.10 (i), we conclude that $\dim\rvert D\rvert>0.$ From the formula (3.2), the arithmetic genus $p_{a}(D)=\frac{1}{2}(4-2+1)(4+2+2)=12.$ Let $E:=K_{S}+D$. The Lemma 2.11 gives that $E$ is very ample on $S$. Thus $E\rvert_{C}$ is very ample on $C$. Thanks to $\deg_{C}E=C\cdot E=4>2g(C)-2=2$, we have that
$h^{0}(C,\mathcal{O}_{C}(E))=\chi(C,\mathcal{O}_{C}(E))=\deg_{C}E+1-p_{a}(C)=4+1-2=3.$
This implies that $C$ is a plane curve with degree $d=C\cdot E=4$. Thus we conclude that the arithmetic genus $p_{\alpha}(C)=\frac{(d-1)(d-2)}{2}=3$. This contradicts that
$p_{\alpha}(C)=\frac{1}{2}(1-1+1)(1+1+2)=2.$
Therefore, the result (i) is true.
(ii) Let curve $C\sim 2H-F$. The formula (3.2) gives that the arithmetic genus
$p_{a}(C)=\frac{1}{2}(2-1+1)(2+1+2)=5.$
According to Theorem 3.7 and the result of (i), we know that $C$ is reduced and irreducible. Assume that $E:=K_{S}+C$. If $h^{0}(S,\mathcal{O}_{S}(C))\geq 2$, from Lemma 2.11, $E$ is very ample on $S$.
From the exact sequence
$0\to\mathcal{O}_{S}(K_{S})\to\mathcal{O}_{S}(E)\to\mathcal{O}_{C}(K_{C})\to 0$
we have that
$h^{0}(S,\mathcal{O}_{S}(E))=h^{0}(C,\mathcal{O}_{C}(K_{C}))=p_{a}(C)=5.$
This shows that $S$ is embedded in $\mathbb{P}^{4}$. This contradicts Theorem 1.3.
Let curve $C\sim H$. The formula (3.2) gives that the arithmetic genus $p_{a}(C)=\frac{1}{2}(1+1)(1+2)=3.$ It is the same as mentioned above, if $h^{0}(S,\mathcal{O}_{S}(C))\geq 2$, we have that $E=K_{S}+C$ is very ample on $S$ and
$h^{0}(S,\mathcal{O}_{S}(E))=h^{0}(C,\mathcal{O}_{C}(K_{C}))=g(C)=3.$
This gives that $S$ is isomorphic to $\mathbb{P}^{2}$, a contradiction. ∎
Proof of Theorem 1.4 of odd type..
(i) Note that if $S$ is a Mori dream space, then $S$ is a fibration with fibers $\mathbb{F}\sim m(H+F)$ or $\mathbb{F}\sim m(H-F)$ for some $m\mathbb{Z}_{>0}.$ The same reason as the proof of Theorem 1.4 of even type, we can get the result.
(ii) By the inequality $K_{S}.C\leq\Delta+2g(C)-2-\chi(C)$, together with $K_{S}.C>0$, we have that the geometric genus $p_{g}(C)\geq 2.$ If $p_{g}(C)=2$, then $K_{S}.C\leq 2.$ This shows that the curve $C\sim H-F$. The Lemma 3.15 provides that $g_{a}(C)=p_{a}(C)=2$. This contradicts Lemma 3.16.
(iii) According to Theorem 3.7 (1), we only consider the curve
$C\sim xH+yF\quad\text{ with}\quad x\geq\rvert y\rvert.$
By the arithmetic genus $p_{a}(C)=\frac{1}{2}(x+y+1)(x-y+2)$, we obtain that
  1. (i)
    if $p_{a}(C)=3$, then $C\sim 2(H-F)$ or $C\sim H+F$ or $C\sim H$;
  2. (ii)
    if $p_{a}(C)=4$, then $C\sim 3(H-F)$;
  3. (iii)
    if $p_{a}(C)=5$, then $C\sim 4(H-F)$ or $C\sim 2(H+F)$ or $C\sim 2H-F$.
Recall that $S$ is not a fibration. This gives that
$h^{0}(S,\mathcal{O}_{S}(mD))\leq 1\quad\text{for}\quad m\in\mathbb{Z}_{>0},\quad D\sim H+F\quad\text{ or}\quad D\sim H-F.$
Together with Lemma 3.16, the proof is complete. ∎