### v10

Abstract.
In this paper, we give a criterion to assess the effectiveness and ampleness of divisors on a fake quadric surface $S$, and then we establish a relationship between the cones:
$\mathring{\mathrm{Eff}}(S)=\mathrm{Amp}(S)\subset\mathrm{SAmp}(S)=\mathrm{Mov}(S)\subset\mathrm{Nef}(S)=\mathrm{Eff}(S)=\overline{\mathrm{Amp}(S)}.$
In particular, we prove that any fake quadric of odd type does not contain a negative curve. This result is central to our manuscript.
As applications, first we give that any fake quadric is a fibration over $\mathbb{P}^{1};$ Subsequently, we show that no fake quadric can be embedded in $\mathbb{P}^{4}$; Finally, we prove that the fake quadric $S$ possesses the bounded cohomology property. This property is characterized by the existence of a positive constant $c_{S}$ such that
$h^{1}(\mathcal{O}_{S}(C))\leq c_{S}h^{0}(\mathcal{O}_{S}(C))$
for any curve $C\subset S$.
1. introduction
The study of fake quadrics is an important topic of algebraic surfaces. A fake quadric $S$ is a minimal projective surface that has the same Betti numbers as the quadric surface but is not biholomorphic to one, or equivalently a minimal surface of general type with numerical invariants:
$q(S)=p_{g}(S)=0,K_{S}^{2}=8,$
where $q(S)$ is the irregularity, $p_{g}(S)$ is the geometric genus and $K_{S}$ is the canonical divisor of $S$ (c.f.[5],[6],[7],[8]).
Fake quadrics are motivated by the study of fake projective planes, that is, surfaces of general type with the same Betti numbers as $\mathbb{P}^{2}$. It is proved by Yau that every fake projective plane is a quotient $\mathbb{B}_{2}/\Gamma$ of the two-dimensional complex ball $\mathbb{B}_{2}$ by a cocompact and torsion-free discrete group $\Gamma$ of Aut$(\mathbb{B}_{2})$ (c.f.[23],[24]). Based on the Yau’s Theorem, a classification of all fake projective planes is done by Prasad and Yeung (c.f.[19]). However, the classification of fake quadrics remains far from being complete. In particular, the existence of a simply connected fake quadric still remains an open question, one that was first raised by Hirzebruch (cf. [6, p.8][12, p.780]).
By the exponential exact sequence, the first Chern class induces an isomorphism $\mathrm{Pic}(S)\cong H^{2}(S,\mathbb{Z})$ for the fake quadrics. Hence the Néron–Severi lattice $\mathrm{N}^{1}(S)\cong H^{2}(S,\mathbb{Z})/\mathrm{tors}$. Furthermore, the Poincaré duality and Hodge index theorem imply that $\mathrm{N}^{1}(S)$ is a unimodular indefinite rank $2$ lattice of signature 00. As a consequence, the Néron–Severi lattice must be even or odd, which means there exist only two cases:
$\mathrm{N}^{1}(S)=U:=\left(\begin{array}[]{cc}0&1\\ 1&0\\ \end{array}\right)\;\;\textrm{ or }\;\;\mathrm{N}^{1}(S)=\langle 1\rangle\oplus\langle-1\rangle:=\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$
For convenience, by a fake quadric of even type we mean its Néron–Severi lattice is an even lattice (i.e., the hyperbolic plane $U$), and by a fake quadric of odd type we mean its Néron–Severi lattice is an odd lattice (i.e., $\langle 1\rangle\oplus\langle-1\rangle$).
There have been many excellent works on fake quadrics (c.f. [7],[8],[10], [16]). In this paper, we discuss the properties of divisors on fake quadrics, a topic that has not been precisely addressed in previous research. Specifically,
  1. (1)
    we establish a criterion for evaluating the effectiveness of divisors;
  2. (2)
    we discuss the sufficient and necessary conditions for determining the ample divisors.
We will investigate these two problems for a fake quadric of even type in Theorem 2.2 (Sec.2) and for a fake quadric of odd type in Theorem 3.7(Sec.3). We believe that this result is new, particularly in those of odd type.
In previous work on the fake quadrics, the odd type was rarely involved. We believe that the possibility of the odd type cannot be ruled out by using the current methods and tools. This led us to have a detailed discussion on this issue. In particular, we provide a detailed argument in Sect.3 to demonstrate that any fake quadric of odd type does not contain negative curve (See Theorem 3.6). This is the key result of this manuscript.
From Theorem 2.2 and Theorem 3.7, the following result can be immediately derived.
Theorem 1.1.
Let $S$ be a fake quadric. Then a relationship between the cones is
$\mathring{\mathrm{Eff}}(S)=\mathrm{Amp}(S)\subset\mathrm{SAmp}(S)=\mathrm{Mov}(S)\subset\mathrm{Nef}(S)=\mathrm{Eff}(S)=\mathbb{R}_{\geq 0}[D_{1}]+\mathbb{R}_{\geq 0}[D_{2}]=\overline{\mathrm{Amp}(S)},$
where $D_{1},D_{2}\subset S$ are two distinct divisors with $D_{i}^{2}=0,i=1,2,$ and $\mathring{\mathrm{Eff}}(S)$ is the set of interior points of $\mathrm{Eff}(S).$
If $S$ is a Mori dream space (The definition of Mori dream space see [10, 13]), by [13, Theorem 2.9], together with Theorem 1.1, we have that
Corollary 1.2.
Let $S$ be a fake quadric. With the above notation, $S$ is a Mori dream space if and only if
$\mathrm{SAmp}(S)=\mathrm{Eff}(S).$
According to Theorem 1.1, we conclude that
Theorem 1.3.
Any fake quadric is a fibration over $\mathbb{P}^{1}.$
Note that the quadric $\mathbb{P}^{1}\times\mathbb{P}^{1}$ is a hypersurface in $\mathbb{P}^{3}$. Furthermore the surface $\mathbb{P}^{2}\sharp\overline{\mathbb{P}^{2}}$ ($\mathbb{P}^{2}$ blowing up along a point) is a subvariety of $\mathbb{P}^{4}$. From this point of view, it is a natural way to ask that is there a fake quadric that can be smoothly embedded in $\mathbb{P}^{4}$? By applying the criterion for divisors, we can derive the following result.
Theorem 1.4.
There is no fake quadric that can be holomorphically embedded in $\mathbb{P}^{4}$.
As a Corollary to this Theorem, we have that
Corollary 1.5.
Let $S$ be a quadric of even type (odd type). For all ample divisors $L\subset S$, if $K_{S}+kL$ is very ample, then $k\geq 2$. Additionally, if $kL$ is very ample, then $k\geq 4$ ($k\geq 3$).
Recall that a smooth projective surface $S$ satisfies bounded cohomology property if there is a positive constant $c_{S}$ such that
$h^{1}(\mathcal{O}_{S}(C))\leq c_{S}h^{0}(\mathcal{O}_{S}(C))$
for every curve $C$ on $S$ (Here the curve $C$ means reduced and irreducible). It arose out of discussions between Harbourne, Roé, Cilberto and Miranda (cf.[3]). As an application of the criterion for divisors, we consider the fake quadrics and have the following Theorem.
Theorem 1.6.
If $S$ is a fake quadric of even type, then the bounded cohomology property holds. If Mori dream space $S$ is a fake quadric of odd type, then the bounded cohomology property is true.
Finally, while this paper introduces the concept of fake quadrics of odd type, the existence of such case remains an open question that requires further investigation. To date, it is still unclear whether a simply connected fake quadric exists. We believe that some of the results presented in this manuscript offer valuable insights into this question.
Acknowledgement.
We are grateful to Yifei Chen, Cheng-Yong Du, and Song Yang for their helpful communications. Special thanks go to Yifan Chen for his detailed discussions on Lemma 3.3, Lemma 3.5, and Theorem 3.6. We would also like to extend our sincere appreciation to Fabrizio Catanese for his valuable insights on the Néron–Severi lattice of fake quadrics and for his thoughtful advice on this manuscript, shared via email.
Notation and conventions
We work throughout over the complex number field $\mathbb{C}$. By a surface we mean a smooth complex projective surface, and if not specifically stated, a curve on a surface always means any non-zero effective divisor on the surface. By a negative curve we mean a reduced and irreducible curve on surface with negative self-intersection. Moreover, we use the following standard notations:
•   $h^{i}(S,\mathcal{O}_{S}(D))$: $=\dim H^{i}(S,\mathcal{O}_{S}(D))$;
•   $C.D$: the intersection number of two divisors $C$ and $D$;
•   $\iota_{\rvert D\rvert}$: a map $\iota_{\rvert D\rvert}:S\to\mathbb{P}^{\dim\rvert D\rvert}$ with $\mathcal{O}(D)=\iota_{\rvert D\rvert}^{\ast}\mathcal{O}_{\mathbb{P}^{\dim\rvert D\rvert}}(1)$;
•  $C\sim D:$ the two divisors $C,D$ are linear equivalence;
•  $C\equiv D:$ the two divisors $C,D$ are numerical equivalence;
•  $C=_{hom}D:$ the two divisors $C,D$ are homology equivalence;
•   $\mathrm{Nef}(S)$: the nef cone of the surface $S$;
•   $\mathrm{Amp}(S)$: the Ample cone of the surface $S$.
•   $\mathrm{SAmp}(S)$: the semiample cone of the surface $S$;
•   $\mathrm{Mov}(S)$: the movable cone of the surface $S$;
•   $\mathrm{Eff}(S)$: the effective cone of the surface $S$.
It is noteworthy that for a fake quadric $S$, by exponent exact sequence, then
$C\sim D\Leftrightarrow C=_{hom}D\quad\text{and}\quad(C=_{hom}D)/\mathrm{Tors}\Leftrightarrow C\equiv D.$
2. Case I. A fake quadric of even type
Let $S$ be a fake quadric surface. It is well-known that the canonical divisor $K_{S}$ is ample. In fact, for any reduced and irreducible curve $C$ on $S$, since $K_{S}$ is nef, so $K_{S}.C=0$ if and only if $C$ is a $(-2)$-curve. Since the number $k$ of $(-2)$-curve is at most $\rho(S)-1=1$ (c.f. [4, VII. Proposition 2.5]), and $k\leq 2(3c_{2}-c_{1}^{2})/9=4/9<1$ (c.f [18, Proposition 2.1.1]), thus $k=0$ which implies $K_{S}$ is ample.
Let $S$ be a fake quadric of even type. By the definition, its Néron–Severi lattice
$\mathrm{N}^{1}(S)=\left(\begin{array}[]{cc}0&1\\ 1&0\\ \end{array}\right)$
and hence there exist two divisors $H$ and $F$ such that
$H^{2}=0,F^{2}=0\textrm{ and }H.F=1.$
From now on, we shall fix two generators $H$ and $F$ in the Néron–Severi lattice $\mathrm{N}^{1}(S)$ of a fake quadric $S$ of even type. Without loss of generality, since $K_{S}^{2}=8$, we may suppose $K_{S}\equiv 2H+2F$. For a divisor $D\equiv xH+yF$, by the Riemann–Roch theorem, we have
$\chi(\mathcal{O}_{S}(D))=1+\frac{1}{2}D(D-K_{S})=(x-1)(y-1).$ (2.1)
2.1. Criterion of divisors on fake quadrics of even type: effectiveness and ampleness
Lemma 2.1.
Let $S$ be a fake quadric of even type and $C$ a reduced and irreducible curve on $S$. If $C\equiv xH+yF$ for some $x,y\in\mathbb{Z}$, then either $x\geq 0$ and $y>0$, or $x>0$ and $y\geq 0$.
Proof.
Since $K_{S}$ is ample, so $K_{S}.C>0$ and thus $x+y>0$. Using the arithmetic genus formula, we have
$p_{a}(C)=\frac{1}{2}(K_{S}.C+C^{2})+1=(x+1)(y+1)\geq 0.$ (2.2)
As a consequence, there hold only two possible cases:
  1. (1)
    $x=-1$ and $y\geq 2$, or $x\geq 2$ and $y=-1$, or
  2. (2)
    $x\geq 0$ and $y>0$, or $x>0$ and $y\geq 0$.
Next we shall rule out the case $(1)$. Suppose $C\equiv-H+yF$ with $y\geq 2$ (or $C\equiv xH-F$ with $x\geq 2$). Then its arithmetic genus $p_{a}(C)=0$ and thus $C$ is a smooth rational curve. Since $K_{S}$ is ample, then $mK_{S}$ is effective for $m\gg 0$ and hence $m(K_{S}+C)$ is effective for $m\gg 0$. By [15, Theorem 1 (2)], we have
$K_{S}.C\leq\Delta+2g(C)-2-\chi(C)=4+0-2-2=0,$
where $\Delta:=3c_{2}-c_{1}^{2}$ and $\chi(C)$ is the topological Euler characteristic of $C$. This contradicts to the ampleness of $K_{S}$ and thus finishes the proof of Lemma 2.1. ∎
Based on the above theorem, we obtain the following result.
Theorem 2.2.
Let $S$ be a fake quadric of even type and $D$ a divisor on $S$. Suppose $D\equiv xH+yF$ for some $x,y\in\mathbb{Z}$.
  1. (i)
    then the cone $\mathrm{Nef}(S)=\mathrm{Eff}(S)=\mathbb{R}_{\geq 0}[H]+\mathbb{R}_{\geq 0}[F].$
  2. (ii)
    the divisor $D$ is ample if and only if $x>0$ and $y>0$.
Proof.
(i) Note that [1]
$\mathrm{SAmp}(S)\subset\mathrm{Mov}(S)\subset\mathrm{Nef}(S)\subset\overline{\mathrm{Eff}(S)}.$
By Lemma 2.1, if $D$ is effective, then $x\geq 0$ and $y\geq 0$. This implies that
$\mathrm{Eff}(S)\subset\mathbb{R}_{\geq 0}[H]+\mathbb{R}_{\geq 0}[F]$
Since
$H^{2}=F^{2}=0\quad\text{and}\quad H.D\geq 0,\quad F.D\geq 0$
for $x\geq 0$ and $y\geq 0$, then $H$ and $F$ are nef divisors. By [14, Theorem 1.5.28], we have that the rays $\mathbb{R}_{\geq 0}[H]$ and $\mathbb{R}_{\geq 0}[F]$ are lie in effective cone. This gives that
$\mathbb{R}_{\geq 0}[H]+\mathbb{R}_{\geq 0}[F]\subset\mathrm{Eff}(S).$
Thus we get that
$\mathrm{Eff}(S)=\mathbb{R}_{\geq 0}[H]+\mathbb{R}_{\geq 0}[F].$
By $\mathrm{Nef}(S)\subset\mathrm{Eff}(S)$, together with the rays $\mathbb{R}_{\geq 0}[H]$ and $\mathbb{R}_{\geq 0}[F]$ lies in $\mathrm{Nef}(S)$, we have that
$\mathbb{R}_{\geq 0}[H]+\mathbb{R}_{\geq 0}[F]\subset\mathrm{Nef}(S).$
Thus the result holds.
(ii) Suppose that $D$ is ample. Then $D^{2}=2xy>0$ and $mD$ is effective for $m\gg 0$. As a result, it follows from (i) that $x>0$ and $y>0$. Conversely, assume $x>0$ and $y>0$ and hence $D^{2}=2xy>0$. For any irreducible curve $C\equiv aH+bF$, thanks to Lemma 2.1, there holds either $a>0,b\geq 0$ or $a\geq 0,b>0$, and thus $D.C=ay+bx>0$. Consequently, the Nakai–Moishezon criterion implies that $D$ is an ample divisor. ∎
According to Theorem 2.2, the Theorem 1.1 for even type is ture.
From the Theorem 2.2 (i) and the formula (2.2), we can immediately obtain that
Corollary 2.3.
Let $S$ be a fake quadric of even type. Then for any curve $C$ on $S$, the $C^{2}\geq 0$ and the arithmetic genus $p_{a}(C)\geq 2$.
As a consequence of the Theorem above, we obtain the following result.
Corollary 2.4.
Let $S$ be a fake quadric of even type and $D$ a divisor on $S$. Suppose $D\equiv xH+yF$ for some $x,y\in\mathbb{Z}$.
  1. (i)
    If $x,y\geq 2$, then $h^{0}(S,\mathcal{O}_{S}(D))>0,$ unless $D\sim K_{S}$.
  2. (ii)
    If $x,y\geq 3$, then $h^{1}(S,\mathcal{O}_{S}(D))=h^{2}(S,\mathcal{O}_{S}(D))=0$ and $h^{0}(S,\mathcal{O}_{S}(D))\geq 4.$
Proof.
(i) Since $K_{S}-D\equiv(2-x)H+(2-y)F$, the assumption and Theorem 2.2 (i) imply that the divisor $K_{S}-D$ is not effective. The Serre duality yields
$h^{2}(S,\mathcal{O}_{S}(D))=h^{0}(S,\mathcal{O}_{S}(K_{S}-D))=0.$
Indeed, if $D\equiv K_{S}$ but $D\not\sim K_{S}$, the $K_{S}-D$ is a torsion divisor and so $h^{0}(S,\mathcal{O}_{S}(K_{S}-D))=0.$ As a result, the Riemann–Roch theorem concludes
$h^{0}(S,\mathcal{O}_{S}(D))\geq\chi(\mathcal{O}_{S}(D))=(x-1)(y-1)\geq 1.$
(ii) By Theorem 2.2 (ii), the $x,y\geq 3$ guarantees that the divisor $D-K_{S}$ is ample. According to the Kodaira vanishing theorem, we have $h^{i}(S,\mathcal{O}_{S}(D))=0$ for $i>0$. Then we conclude that
$h^{0}(S,\mathcal{O}_{S}(D))=\chi(S,\mathcal{O}_{S}(D))=(x-1)(y-1)\geq 4.$
This completes the proof. ∎
2.2. Any fake quadric of even type is a fibration.
In this section, we shall prove that the fake quadric of even type is a fibration. Before the proof of main result, some useful Lemmas are needed.
Lemma 2.5.
If $\pi:S\to C$ is a fibration with irregularity $q(S)=0$, then the genus $g(C)=0$ and $C\cong\mathbb{P}^{1}$.
Proof.
Assume that $g(C)\neq 0$. Let $\alpha\in H^{1,0}(C)$ be a nonzero element. Then we have that $\pi^{\ast}\alpha\neq 0$. This shows that $H^{1,0}(S)\neq 0$. This contradicts the irregularity $q(S)=0$. ∎
Lemma 2.6.
[1, Proposition 1.1] Let $X$ be a smooth projective surface with $\rho(X)\geq 3$. Suppose that the $\mathrm{Eff}(X)$ is polyhedral. Then
$\mathrm{Eff}(X)=\sum_{[C]\in\mathrm{Exc}[X]}\mathbb{R}_{\geq 0}[C]$
where $\mathrm{Exc}[X]$ is the set of classes of negative curves $C$ on $X$.
Now, we begin to prove the Theorem
Proof of Theorem 1.3 of even type..
Suppose that the fake quadric $S$ is not a fibration. Then
$h^{0}(S,\mathcal{O}_{S}(kH))\leq 1,\quad h^{0}(S,\mathcal{O}_{S}(kF))\leq 1\quad\text{for}\quad\text{ any}\quad k>0.$
Assume that $D_{1}\in\mathbb{R}_{\geq 0}[H]$ and $D_{2}\in\mathbb{R}_{\geq 0}[F]$ are reduced and irreducible curves. Let point $p\in D_{1},p\notin D_{2}$ and $p$ is not a singular point of $D_{1}$. Blowing up $S$ along $p$, $\pi:\tilde{S}\to S$, we have that
$\mathbb{R}_{\geq 0}[E]+\mathbb{R}_{\geq 0}[D_{1}-E]+\mathbb{R}_{\geq 0}[D_{2}]\subset\mathrm{Eff}(\tilde{S}),$
where $E$ is the exception divisor. Without loss of generality, we denote $\pi^{\ast}D_{i},i=1,2$ as $D_{i}$ (See figure 1).
Follows by Lemma 2.6, we have that there exists a ray
$\mathbb{R}_{\geq 0}[D_{2}-mE]\subset\mathrm{Eff}(\tilde{S}),\quad m\in\mathbb{Q}_{>0}.$
Otherwise, $\mathbb{R}_{\geq 0}[D_{2}]$ is an extramel ray. According to Lemma 2.6, $\mathrm{Eff}(\tilde{S})$ is not polyhedral. Since $\mathbb{R}_{\geq 0}[E]$ and $[D_{1}-E]$ are the extremal rays, this gives that the ray $\mathbb{R}_{\geq 0}[D_{2}]$ is in the set of interior points of $\mathrm{Eff}(\tilde{S}),$ a contradiction.
This shows that there exists a curve $C\in\mathbb{R}_{\geq 0}[D_{2}-mE]$. Since $C+mE\in|D_{2}|$, then we get that $h^{0}(S,\mathcal{O}_{S}(kD_{2}))\geq 2$ for $k\gg 0$. This contradicts that $h^{0}(S,\mathcal{O}_{S}(kD_{2}))\leq 1$ for any $k>0$. ∎
$\mathbb{R}_{\geq 0}[D_{1}]$$\mathbb{R}_{\geq 0}[D_{2}]$$\mathbb{R}_{\geq 0}[E]$$\mathbb{R}_{\geq 0}[D_{1}-E]$$\mathbb{R}_{\geq 0}[D_{2}-mE]$
Figure 1. Effective cones of $\tilde{S}$.
2.3. Embedding fake quadric of even type in $\mathbb{P}^{4}$
It is well-known that every smooth projective surface is holomorphically embedded in $\mathbb{P}^{5}$. However, some of smooth projective surfaces can be embedded in $\mathbb{P}^{4}$, such as the projective plane, the quadric. As a by-product, we have an observation (Lemma 1.4) which we believe is new.
Proof of Theorem 1.4 of even type..
Let $S$ be a fake quadric and $\iota:S\hookrightarrow\mathbb{P}^{4}$ an embedding. Then there exists a very ample divisor $D$ such that $\mathcal{O}_{S}(D)=\iota^{\ast}\mathcal{O}_{\mathbb{P}^{4}}(1)$. Moreover, the fake quadric $S$ satisfies the following numerical condition (c.f. [11, Appendix A, Example 4.1.3]):
$d^{2}-10d-5D.K_{S}-2K_{S}^{2}+12+12p_{a}(S)=0,$ (2.3)
where $d=D^{2}$ and the arithmetic genus $p_{a}(S)=0$. Assume that $D\equiv xH+yF$. It follows from Theorem 2.2 that $x>0$ and $y>0$. Therefore, the equation (2.3) implies that $x$ and $y$ must satisfy
$02x^{2}y^{2}-10xy-5x-5y-2=0,\quad\text{for}\quad x>0,y>$ 0 (2.4)
By [25, Chapter V, Corollary 2.9], we have that $h^{0}(S,\mathcal{O}_{S}(D))<\binom{4}{2}=6$. Together with Corollary 2.4 (ii), we have that
$(x-1)(y-1)=h^{0}(S,\mathcal{O}_{S}(D))<6,\quad\text{for}\quad x\geq 3,y\geq 3.$ (2.5)
If $30<x<$ 3 or $30<y<$ 3, there is no solution of equation (2.4). For $x\geq 3,y\geq 3$, combine with (2.4) and (2.5), it can be concluded that the values of $x,y$ do not exist. This completes the proof. ∎
Now, we start to prove the Corollary 1.5 of even type.
Proof of Corollary 1.5 of even type..
According to Theorem 2.2, the divisor $L\equiv H+F$ is ample. Conisder the divisor $K_{S}+L$. Follows by Corollary 2.4, we have
$h^{0}(S,\mathcal{O}_{S}(K_{S}+L))=\chi(\mathcal{O}_{S}(K_{S}+L))=\chi(\mathcal{O}_{S}(3H+3F))=(3-1)(3-1)=4.$
If such divsior $K_{S}+L$ is very ample, then $\tau_{\rvert K_{S}+L\rvert}:S\to\mathbb{P}^{3}$ is an embedding. This contradicts the Theorem 1.4.
Since $3L\sim 3H+3F\sim K_{S}+L,$ for a similar reason as described above, the $3L$ is not very ample. ∎
2.4. The bounded cohomology property for the fake quadrics of even type
In this subsection, a curve means a reduced and irreducible curve on a surface. Before the main result of this subsection, the following Lemmas are necessary.
Lemma 2.7.
If $S$ is a fake quadric of even type and a curve $C\equiv xH+yF,x,y\in\mathbb{Z}$ on $S$, then $C$ satisfies the bounded cohomology property for $x,y\geq 2$ or $x=1$ or $y=1$.
Proof.
Recall that the divisor $K_{S}-C$ is not effective. The Serre duality yields
$h^{2}(S,\mathcal{O}_{S}(C))=h^{0}(S,\mathcal{O}_{S}(K_{S}-C))=0.$
As a result, the Riemann–Roch theorem concludes
$h^{0}(S,\mathcal{O}_{S}(C))-h^{1}(S,\mathcal{O}_{S}(C))=\chi(\mathcal{O}_{S}(C))=(x-1)(y-1)$
If $x,y\geq 2$, then we have that $h^{1}(S,\mathcal{O}_{S}(C))<h^{0}(S,\mathcal{O}_{S}(C)).$ If $x=1$ or $y=1$, we have that $h^{1}(S,\mathcal{O}_{S}(C))=h^{0}(S,\mathcal{O}_{S}(C)).$ To summary, we have that
$h^{1}(S,\mathcal{O}_{S}(C))\leq h^{0}(S,\mathcal{O}_{S}(C)).$
Now we begin to prove the Theorem 1.6
Proof of Theorem 1.6 of even type..
Under the above discussion, we only consider that $x=0$ or $y=0$. Without loss of generality, we suppose that $C\equiv x_{0}H$ for some $x_{0}>0.$ Since $C$ is reduced and irreducible, the $C^{2}=0$ gives that $h^{0}(S,\mathcal{O}_{S}(C))$ is either $1$ or $2$.
If $h^{0}(S,\mathcal{O}_{S}(C))=1$, then there are finitely many $x_{0}$. Let $c_{1}:=max\{x_{0}\}.$ According to
$h^{0}(S,\mathcal{O}_{S}(C))-h^{1}(S,\mathcal{O}_{S}(C))=\chi(\mathcal{O}_{S}(C))=(x-1)(y-1)=-(x_{0}-1),$
togethr with $h^{0}(S,\mathcal{O}_{S}(C))=1$, we have that $h^{1}(S,\mathcal{O}_{S}(C))=x_{0}$. Then we have that
$h^{1}(S,\mathcal{O}_{S}(C))=x_{0}h^{0}(S,\mathcal{O}_{S}(C))$
and so
$h^{1}(S,\mathcal{O}_{S}(C))\leq c_{1}h^{0}(S,\mathcal{O}_{S}(C)).$
If $h^{0}(S,\mathcal{O}_{S}(C))=2$, the same reason as above, we have that $h^{1}(S,\mathcal{O}_{S}(C))=x_{0}+1$ and so $h^{1}(S,\mathcal{O}_{S}(C))=\frac{1}{2}(x_{0}+1)h^{0}(S,\mathcal{O}_{S}(C))$. In this case, the the number of $x_{0}$ is also finite. This shows that
$h^{1}(S,\mathcal{O}_{S}(C))\leq c_{2}h^{0}(S,\mathcal{O}_{S}(C))$
for $c_{2}:=max\{\frac{1}{2}(x_{0}+1)\}.$ This implies that the bounded cohomology property holds. ∎
3. Case II. A fake quadric of odd type
Let $S$ be a fake quadric of odd type. According to the definition, the Néron–Severi lattice
$\mathrm{N}^{1}(S)=\left(\begin{array}[]{cc}1&0\\ 0&-1\\ \end{array}\right)$
and thus there are two divisors $H$ and $F$ with intersection numbers
$H^{2}=1,F^{2}=-1\textrm{ and }H.F=0.$
From now on, we will fix two generators $H$ and $F$ of the Néron–Severi lattice $\mathrm{N}^{1}(S)$. Note that the canonical divisor $K_{S}$ is a primitive divisor in $\mathrm{N}^{1}(S)$ (c.f. [21, Theorems 3.1 (iii), 3.10]). Without loss of generality, since $K_{S}^{2}=8$, we may assume $K_{S}\equiv 3H-F$. For a divisor $D\equiv xH+yF$, by the Riemann–Roch theorem, we have
$\chi(\mathcal{O}_{S}(D))=\frac{1}{2}(x+y-1)(x-y-2).$ (3.1)
3.1. Criterion of divisors on fake quadrics of odd type: effectiveness and ampleness
Lemma 3.1.
Let $S$ be a fake quadric of odd type and $C$ a reduced and irreducible curve on $S$. If $C\equiv xH+yF$ for some $x,y\in\mathbb{Z}$, then $x\geq 0$ and $-x-1<y\leq x+1$.
Proof.
Since $K_{S}$ is ample, then $K_{S}.C>0$ and thus $03x+y>$ 0. By the arithmetic genus formula, we get
$p_{a}(C)=\frac{1}{2}(x+y+1)(x-y+2)\geq 0.$ (3.2)
Then we have only two possible cases:
  1. (1)
    $x=0$ and $y=2$, or $x>0$ and $y=-(x+1)$, or $x>0$ and $y=x+2$, or
  2. (2)
    $x\geq 0$ and $-(x+1)<y\leq x+1$.
We will rule out the case $(1)$. The idea is as same as the proof of Lemma 2.1. Suppose $C\equiv 2F$ or $C\equiv xH-(x+1)F$ or $C\equiv xH+(x+2)F$ with $x>0$. Then the arithmetic genus $p_{a}(C)=0$ and thus $C$ is a smooth rational curve. Since $K_{S}$ is ample, so $mK_{S}$ is effective for $m\gg 0$ and thus $m(K_{S}+C)$ is effective for some $m\gg 0$. By [15, Theorem 1 (2)], we have $K_{S}.C\leq\Delta+2g(C)-2-\chi(C)=4+0-2-2=0$. This contradicts to the ampleness of $K_{S}$. ∎
To more accurately characterize the divisors, in the following we mainly focus on the existence of negative curves on fake quadrics. (Here the negative curve means reduced and irreducible).
Lemma 3.2.
If the Néron–Severi lattice of fake quadric $S$ is odd type, then there is at most one negative curve on $S$.
Proof.
According to Lemma 3.1, any negative curve (reduced and irreducible) is numerically equivalent to $xH+(x+1)F$ for some $x\in\mathbb{Z}_{\geq 0}$. Suppose that $x_{0}\in\mathbb{Z}_{\geq 0}$ is the minimal number such that $x_{0}H+(x_{0}+1)F$ is numerically equivalent to an effective divisor. Let curve $E\equiv x_{0}H+(x_{0}+1)F$. The $E^{2}<0$ gives that the linear system $\rvert E\rvert$ contains a fixed component. This implies that $\dim\rvert E\rvert=0.$
For any divisor $D\equiv xH+(x+1)F,x>x_{0}$, we may conclude that if $D$ effective, then all ellements of the linear system $\rvert D\rvert$ are not reduced or irreducible. Otherwise, then $\dim\rvert D\rvert=0$ and so $D-E$ is not effective. By $\deg D\rvert_{E}<0,$ we have $h^{0}(E,\mathcal{O}_{E}(D))=0.$ According to the exact sequence
$0\to h^{0}(S,\mathcal{O}_{S}(D-E))\to h^{0}(S,\mathcal{O}_{S}(D))\to h^{0}(E,\mathcal{O}_{E}(D))\to\ldots$
together with $h^{0}(S,\mathcal{O}_{S}(D-E))=0,$ we have $h^{0}(S,\mathcal{O}_{S}(D))=0$, a contradiction. Thus any negative curve is numerically equivalent to $E$.
Suppose that $E^{\prime}\equiv E$ is another negative curve. Then $E$ and $E^{\prime}$ only differ by a torsion divisor. This shows that there exists a $m\in\mathbb{Z}_{>0}$ such that $mE\sim mE^{\prime}$. Thus we that $h^{0}(S,\mathscr{O}_{S}(mE))\geq 2$. Since $E^{2}<0$, by the exact sequence
$0\to h^{0}(S,\mathcal{O}_{S}(kE-E))\to h^{0}(S,\mathcal{O}_{S}(kE))\to h^{0}(E,\mathcal{O}_{E}(kE))\to\ldots,$
then
$h^{0}(S,\mathscr{O}_{S}(mE))=h^{0}(S,\mathcal{O}_{S}(mE-E))=\cdots=h^{0}(S,\mathcal{O}_{S}(E))=1,$
a contradiction. ∎
Lemma 3.3.
If the Néron–Severi lattice of fake quadric $S$ is odd type and the negative curve $E\equiv x_{0}H+(x_{0}+1)F$ is existed, then
  1. (i)
    the cones satisfy
    $\mathrm{Eff}(S)=\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[E]\quad\text{and}\quad\mathrm{Nef}(S)=\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[N],$
    where $N\equiv(x_{0}+1)H+x_{0}F$;
  2. (ii)
    if any divisor
    $Z\in\mathbb{R}_{\geq 0}[N]+\mathbb{R}_{\geq 0}[E],\quad Z\notin\mathbb{R}_{\geq 0}[N],$
    then the linear system $|Z|$ has a fixed component and the support of such fixed component is $E$;
  3. (iii)
    the divisor $N$ is semiapmle.
[…]
Figure 2. Cones of a fake quadric of odd type
Proof.
(i) It is shown in Lemma 3.1 and Lemma 3.2 that
$\mathrm{Nef}(S)\subseteq\overline{\mathrm{Eff}(S)}\subseteq\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[E].$
If the divisor
$D\in\mathbb{R}_{\geq 0}[N]+\mathbb{R}_{\geq 0}[E]$
then we have $D.E\leq 0$. This gives that the cone $\mathbb{R}_{\geq 0}[E]+\mathbb{R}_{\geq 0}[N]$ does not contain any ample divisor. By $\overline{\mathrm{Amp}(S)}=\mathrm{Nef}(S),$ we have that
$\mathrm{Nef}(S)\subseteq\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[N].$
Since any divisor $C\in\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[E]$ satisfies that $C.(H-F)\geq 0$, together with the self-number $(H-F)^{2}=0$, we conclude that $H-F$ is a nef divisor. The same argument as stated above, we can also prove that $N$ is a nef divisor. Thus we get that $\mathrm{Nef}(S)=\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[N].$ Follows by $\overline{\mathrm{Eff}(S)}$ is dual to $\mathrm{Nef}(S)$, we conclude that
$\overline{\mathrm{Eff}(S)}=\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[E].$
From [14, Theorem 1.5.28], we have that the divisor $A:\equiv m(H-F)$ is effective for $m$ sufficiently large. Therefore, the ray $\mathbb{R}_{\geq 0}[H-F]\subset\mathrm{Eff}(S).$ Together with $\mathbb{R}_{\geq 0}[E]\subset\mathrm{Eff}(S),$ one can get that
$\mathrm{Eff}(S)=\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[E].$
(ii) If there exists a divsior
$Z\in\mathbb{R}_{\geq 0}[N]+\mathbb{R}_{\geq 0}[E],\quad Z\notin\mathbb{R}_{\geq 0}[N]$
satisfies that the linear system $|Z|$ does not contain fixed part, then the base locus Bs$|Z|$ is a finite set. It is shown in [14, P.132] that $Z$ is semi-ample. This contradicts that
$\mathrm{SAmp}(S)\subseteq\mathrm{Nef}(S)=\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[N].$
The support of fixed component will be discussed in (iii).
(iii) Note that if divisor $D\in\mathrm{Eff}(S),D\notin\mathbb{R}_{\geq 0}[E],\mathbb{R}_{\geq 0}[H-F]$, then $h^{0}(mD)$ grows in $m^{2}.$ Indeed, if $D^{2}\leq 0$, then $D\in\mathbb{R}_{\geq 0}[N]+\mathbb{R}_{\geq 0}[E]$. By Zariski decomposition, we have that $h^{0}(mD)=h^{0}(mN)$ and so the result holds. This result implies that the stable base locus $B(D):=\cap_{m\geq 1}Bs(|mD|)$ only possibly contains two kind of curves, which are lie in the rays $\mathbb{R}_{\geq 0}[E]$ and $\mathbb{R}_{\geq 0}[H-F]$.
If $N$ is not semiample, then the stable base locus $B(N)\neq\emptyset.$ The [14, P.132] gives that $B(N)$ must contain curves. From Wilson’s Theorem [14, Theorem 2.3.9], we have
$mN\equiv A+B,\quad m\gg 0,$
where $A\in\mathrm{Nef}(S)$ is semiample and $B\in\mathbb{R}_{\geq 0}[N]+\mathbb{R}_{\geq 0}[E]$ is effective. According to (ii), we have that
$nB\equiv B_{1}+B_{2},\quad n\gg 0,$
where $B_{1}\in\mathrm{Nef}(S)$ is semiample and $B_{2}\in\mathbb{R}_{\geq 0}[N]+\mathbb{R}_{\geq 0}[E]$ is fixed part. Then we conclude that the support
$\mathrm{Supp}B_{2}=\mathrm{Supp}B(B)=\mathrm{Supp}B(N)=E.$
Suppose that $mN\equiv A+E_{0}$, where $E_{0}=k_{0}E,k_{0}>0$ is fixed part. This gives that
$m^{2}N^{2}=\mathrm{Vol}(mN)=\mathrm{Vol}(A)=A^{2}.$
From
$A^{2}=m^{2}N^{2}=(A+E_{0})^{2}=A^{2}+2AE_{0}+E_{0}^{2},$
we have that $02AE=-E_{0}^{2}=k_{0}^{2}>$ 0 and so $AE>0$. This shows that
$A\in\mathbb{R}_{\geq 0}[H]+\mathbb{R}_{\geq 0}[H-F]\quad\text{and}\quad A\notin R_{\geq 0}[H]$
Thus we get that $A\equiv m(x_{0}+1)H-l_{0}F$ and $E_{0}\equiv(mx_{0}+l_{0})E$ for some $l_{0}\geq 0.$ Let $D\equiv(x_{0}+1)H$ and consider
$mG\equiv mD+E_{0}\in\mathbb{R}_{\geq 0}[N]+\mathbb{R}_{\geq 0}[E].$
By Zariski decomposition
$mG\equiv mN+l_{0}E,$
we conclde that $\mathrm{Vol}(mG)=\mathrm{Vol}(mN).$
If $x_{0}>0,$ then $D$ is ample. This implies that
$\mathrm{Vol}(mG)\geq\mathrm{Vol}(mD)=m^{2}D^{2}>m^{2}N^{2}=\mathrm{Vol}(mN),$
a contradiction. Thus $N$ is semiample.
If $x_{0}=0,$ then $N\equiv H$ and $E\equiv F.$ Note that $3mN-E\equiv K_{S}+3(n-1)H$. Since $H$ is big and nef, by Kawamata-Viehweg Theorem, we have $h^{i}(S,\mathcal{O}_{S}(3mN-E))=0$ for $i=1,2$. This implies that the trace of linear system $|3mN|$ on $E$ is complete and so semiample. Thus the fixed part of $|3nN|$ does not contain $E$. This contradicts that $\mathrm{Supp}B(N)=E$. Then $N$ is semiample. ∎
The following result is very helpful for the proof of the Lemma 3.5.
Proposition 3.4.
[4, III. (2.1) Theorem] [Grauert’s criterion] A reduced, compact connected curve $C$ with irreducible components $C_{i}$ on a smooth surface is exceptional if and only if the intersection matrix $(C_{i}C_{j})$ is negative definite. Furthermore, a surface obtained by contracting curve $C$ to a point is a normal surface.
Lemma 3.5.
With above condition,
  1. (i)
    if the morphism $f:S\to S^{\prime}$ contracts $E$ to a point, then $S^{\prime}$ is $\mathbb{Q}$-Gorenstein normal projective surface;
  2. (ii)
    the negative curve $E\equiv F$.
Proof.
(i) With above notation, by $N.E=0,$ the birational morphism
$f:=\iota_{\rvert mN\rvert}:S\to S^{\prime}\subset\mathbb{P}^{\rvert mN\rvert}\quad\text{for}\quad\text{some}\quad m>0$
contracts $E$ to a point. Since $C^{2}<0$, according to Grauert’s criterion (Propositon 3.4), we may conclude that $S^{\prime}$ is a normal analytic surface and so a normal projective surface.
Denote that $M:=f_{\ast}H$. By
$f_{\ast}[x_{0}H+(x_{0}+1)F]\equiv f_{\ast}E=0,$
we have that $f_{\ast}F\equiv-\frac{x_{0}}{x_{0}+1}M$. Note that the canonical divisor
$K_{S^{\prime}}=f_{\ast}K_{S}=3f_{\ast}H-f_{\ast}F\equiv 3M+\frac{x_{0}}{x_{0}+1}M=\frac{4x_{0}+3}{x_{0}+1}M.$ (3.3)
Since the Picard number $\rho(S^{\prime})=1$, we can assume that
$\mathcal{O}_{S^{\prime}}(m[(x_{0}+1)H+x_{0}F])\equiv\mathcal{O}_{S^{\prime}}(mN)\equiv f^{\ast}\mathcal{O}_{S^{\prime}}(rM)\equiv f^{\ast}\mathcal{O}_{S^{\prime}}(1),\quad r\in\mathbb{R}.$ (3.4)
By the projection formula, we have that
$\mathcal{O}_{S^{\prime}}(rM)=\mathcal{O}_{S^{\prime}}(1)=f_{\ast}\mathcal{O}_{S}\otimes_{\mathcal{O}_{S^{\prime}}}\mathcal{O}_{S^{\prime}}(1)=f_{\ast}(\mathcal{O}_{S}\otimes_{\mathcal{O}_{S}}f^{\ast}\mathcal{O}_{S^{\prime}}(1))=f_{\ast}f^{\ast}\mathcal{O}_{S^{\prime}}(1)=f_{\ast}f^{\ast}\mathcal{O}_{S^{\prime}}(rM).$
From (3.4), this indicates that
$rM=f_{\ast}f^{\ast}(rM)\equiv$ $mf_{\ast}[(x_{0}+1)H+x_{0}F]=\frac{m(2x_{0}+1)}{x_{0}+1}M$
and then $\frac{m}{r}=\frac{x_{0}+1}{2x_{0}+1}$. From the (3.3), (3.4) and $r\in\mathbb{Q}$, we can get that $K_{S^{\prime}}$ is $\mathbb{Q}$-Cartier divisor and so $S^{\prime}$ is $\mathbb{Q}$-Gorenstein normal variety.
(ii) According to (3.4), we conclude that
$f^{\ast}M\equiv\frac{m}{r}[(x_{0}+1)H+x_{0}F]\equiv\frac{x_{0}+1}{2x_{0}+1}[(x_{0}+1)H+x_{0}F].$
Together with (3.3), we get that
$f^{\ast}K_{S^{\prime}}\equiv\frac{4x_{0}+3}{2x_{0}+1}[(x_{0}+1)H+x_{0}F].$
Follows by
$f^{\ast}K_{S^{\prime}}=K_{S}+kE,\ k\in\mathbb{Q},$
we have that
$\frac{4x_{0}+3}{2x_{0}+1}[(x_{0}+1)H+x_{0}F]\equiv(3H-F)+k[x_{0}H+(x_{0}+1)F]$
and so
$[(4-2k)x_{0}^{2}-(k-1)x_{0}]H+[(4-2k)x_{0}^{2}-(3k-5)x_{0}-(k-1)]F\equiv 0.$
Thus we get the equations
$(4-2k)x_{0}^{2}=(k-1)x_{0},\quad(4-2k)x_{0}^{2}=(3k-5)x_{0}+k-1.$
And then we have that
$x_{0}=0\quad\text{or}\quad-1\quad\text{and}\quad k=1\quad\text{or}\quad 3,$
Since $x_{0}\geq 0$, then $x_{0}=0$ and $k=1.$ This shows that
$E\equiv F\quad\text{and}\quad f^{\ast}K_{S^{\prime}}=K_{S}+E.$
Theorem 3.6.
Let $S$ be a fake quadric of odd type. Then the negative curve on $S$ does not exist.
Proof.
Suppose that the negative curve exist. With the same notations as above, the reduced and irreducible negative curve $E\equiv F.$ For the big and nef divisor
$A:=(k-3)H-(m-1)F,\quad 0\leq m-1<k-3,$
the $K_{S}+A\equiv kH-mF.$ According to Kawamata-Viehweg vanishing Theorem, we have that
$h^{0}(S,\mathcal{O}_{S}(kH-mF))=\chi(S,\mathcal{O}_{S}(kH-mF))$
This gives that
$h^{0}(S,\mathcal{O}_{S}(kH-m_{0}F))=\chi(S,\mathcal{O}_{S}(kH-m_{0}F))$
$=1+\frac{1}{2}((kH-m_{0}F)^{2}+(kH-m_{0}F).(3H-F))=\frac{1}{2}(k^{2}+3k-m_{0}^{2}-m_{0})+1.$
for $k$ sufficiently large and $m_{0}>0$. From the exact sequence,
$0\to\mathcal{O}_{S}(kH-F)\to\mathcal{O}_{S}(kH)\to\mathcal{O}_{F}\to 0,$
we have that
$h^{0}(S,\mathcal{O}_{S}(kH))=h^{0}(S,\mathcal{O}_{S}(kH-F))+1,k\gg 0.$
In the following, we will show that the base locus of the linear system $|kH-F|,k\gg 0$ is a finite set. Assume that linear system $|kH-F|$ has fixed component. Follows from Lemma 3.3 (see Figure 2), the fixed component is only possibility three cases
$C\equiv r_{0}H-F,\quad C\equiv r_{0}H,\quad C\equiv r_{0}F,\quad r_{0}>0.$
Case (i). For the reason as above, by
$h^{0}(S,\mathcal{O}_{S}(kH-F))=\chi(S,\mathcal{O}_{S}(kH-F))>\chi(S,\mathcal{O}_{S}((k-r_{0})H-F))+1$
$=h^{0}(S,\mathcal{O}_{S}((k-r_{0})H-F))+1=h^{0}(S,\mathcal{O}_{S}((k-r_{0})H))=h^{0}(S,\mathcal{O}_{S}(kH-F-C))$
for $k\gg 0,$ we have that $C\equiv r_{0}H-F$ does not the fixed component of $|kH-F|$.
Case (ii). For the reason as above, by
$h^{0}(S,\mathcal{O}_{S}(kH-F))=\chi(S,\mathcal{O}_{S}(kH-F))>\chi(S,\mathcal{O}_{S}((k-r_{0})H-F))$
$=h^{0}(S,\mathcal{O}_{S}((k-r_{0})H-F))=h^{0}(S,\mathcal{O}_{S}(kH-F-C))$
for $k\gg 0,$ we have that $C\equiv r_{0}H$ does not the fixed component of $|kH-F|$.
Case (iii). For the reason as above, by
$h^{0}(S,\mathcal{O}_{S}(kH-F))=\chi(S,\mathcal{O}_{S}(kH-F))>\chi(S,\mathcal{O}_{S}(kH-(r_{0}+1)F))$
$=h^{0}(S,\mathcal{O}_{S}(kH-(r_{0}+1)F))=h^{0}(S,\mathcal{O}_{S}(kH-F-C))$
for $k\gg 0,$ we have that $C\equiv r_{0}F$ does not the fixed component of $|kH-F|$.
From the above reason, the base locus of the linear system $|kH-F|,\ k\gg 0$ is a finite set. Thus we can choose a pencil $L\subset|kH-F|$, which does not contain any fixed part. Note that $kH-2F\equiv K_{S}+[(k-3)H-F].$ Since $(k-3)H-F$ is big and nef, by kawamata-Viehweg Theorem, we have $h^{1}(S,\mathcal{O}_{S}(kH-2F))=0.$ This implies that the trace of linear system $|kH-F|$ on $E$ is complete and so base point free. Thus any irreducible component of each element of $L$ does not contain $E$. This gives that all base points of $L$ are not in $E$. Otherwise, if we blow up such a base point in $E$, the condition $(kH-F).E=1$ implies that the intersection number of the strict tranform of $D\in L$ and $E$ would be negative. This gives that $E$ is contained in an element of $L$, a contradiction.
Blowing up all the base points of $L$, we get a fibration over $\mathbb{P}^{1}.$ From the above reason, the $(kH-F).E=1$ implies that the strict transform of $E$ is a section of such fibration. Since all base points of $L$ are not in $E$, then the strict transform of $E$ is isomorphic to $E$ and so $E\cong\mathbb{P}^{1}.$ This contradicts that the arithmetic genus $p_{a}(E)=p_{a}(F)=1.$
From the above results, it can be inferred that
Theorem 3.7.
Let $S$ be a fake quadric of odd type and $D$ a divisor on $S$. Assume $D\equiv xH+yF$ for $x,y\in\mathbb{Z}$.
  1. (i)
    then the cone $\mathrm{Nef}(S)=\mathrm{Eff}(S)=\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[H+F].$
  2. (ii)
    the divisor $D$ is ample if and only if $x>\rvert y\rvert$
Proof.
(i) It follows immediately from Lemma 3.1 and Lemma 3.5 that if $D$ is effective, then $x\geq\rvert y\rvert$. This implies that
$\mathrm{Eff}(S)\subset\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[H+F].$
Since
$(H-F)^{2}=(H+F)^{2}=0\quad\text{and}\quad(H-F).D\geq 0,\quad(H+F).D\geq 0$
for $x\geq\rvert y\rvert$, then $H-F$ and $H+F$ are nef divisors. By [14, Theorem 1.5.28], we have that the rays $\mathbb{R}_{\geq 0}[H-F]$ and $\mathbb{R}_{\geq 0}[H+F]$ are lie in effective cone. This gives that
$\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[H+F]\subset\mathrm{Eff}(S).$
By $\mathrm{Nef}(S)\subset\mathrm{Eff}(S)$, together with the rays $\mathbb{R}_{\geq 0}[H-F]$ and $\mathbb{R}_{\geq 0}[H+F]$ lies in $\mathrm{Nef}(S)$, we have that
$\mathbb{R}_{\geq 0}[H-F]+\mathbb{R}_{\geq 0}[H+F]\subset\mathrm{Nef}(S).$
Thus the result holds.
(ii) Suppose $D$ is ample. Then $D^{2}=x^{2}-y^{2}>0$ and $mD$ is effective for $m\gg 0$. As a result, it follows from (i) that $x>\rvert y\rvert$. Conversely, assume $x>\rvert y\rvert$ and hence $D^{2}=x^{2}-y^{2}>0$. For any irreducible curve $C\equiv aH+bF$, thanks to (i) ($a\geq\rvert b\rvert$), thus $D.C=ax-by>0$. Consequently, the Nakai–Moishezon criterion implies that $D$ is an ample divisor. ∎
According to Theorem 3.7, the Theorem 1.1 for odd type is ture.
As a by-product, together with the arithmetic genus formula (3.2), we have the following Corollary.
Corollary 3.8.
Let $S$ be a fake quadric of odd type. Then for any curve $C$ on $S$, the $C^{2}\geq 0$ and the arithmetic genus $p_{a}(C)\geq 2$.
As another consequence of Theorem 3.7, we can obtain the following result.
Corollary 3.9.
Let $S$ be a fake quadric of odd type and $D$ a divisor on $S$. Suppose $D\equiv xH+yF$.
  1. (i)
    If $x+y>1$ and $x-y>2$, then $h^{0}(S,\mathcal{O}_{S}(D))>0$ unless $D\sim K_{S}$;
  2. (ii)
    If $x-3>\rvert y+1\rvert$, then $h^{1}(S,\mathcal{O}_{S}(D))=h^{2}(S,\mathcal{O}_{S}(D))=0$ and $h^{0}(S,\mathcal{O}_{S}(D))\geq 3.$
Proof.
(i) Since $K_{S}-D\equiv(3-x)H-(y+1)F$, the assumptions and Theorem 3.7 imply that the divisor $K_{S}-D$ is not effective. Indeed, by $x+y>1$ and $x-y>2$, we have that $x\geq 3$. This gives that $(3-x)<\rvert y+1\rvert$ unless $D\equiv K_{S}$. By the Serre duality and Theorem 3.7 (i), we obtain $h^{2}(S,\mathcal{O}_{S}(D))=h^{0}(S,\mathcal{O}_{S}(K_{S}-D))=0.$ If $D\equiv K_{S}$ but $D\not\sim K_{S}$, the $K_{S}-D$ is a torsion divisor and so $h^{0}(S,\mathcal{O}_{S}(K_{S}-D))=0.$ As a result, the Riemann–Roch theorem implies
$h^{0}(S,\mathcal{O}_{S}(D))\geq\chi(S,\mathcal{O}_{S}(D))=\frac{1}{2}(x+y-1)(x-y-2)\geq 1.$
(ii) By assumption, the divisor $D-K_{S}=(x-3)H+(y+1)F$ satisfies $x-3>\rvert y+1\rvert$. Hence, Theorem 3.7 implies that the divisor $D-K_{S}$ is ample. Moreover, according to the Kodaira vanishing theorem, we have $h^{i}(S,\mathcal{O}_{S}(D))=0$ for $i>0$. The Riemann–Roch theorem yields
$h^{0}(S,\mathcal{O}_{S}(D))=\chi(S,\mathcal{O}_{S}(D))=\frac{1}{2}(x+y-1)(x-y-2)\geq 3.$
This completes the proof. ∎
3.2. Any fake quadric of odd type is a fibration.
Proof of Theorem 1.3 of odd type..
We replace $H,F$ of even type with $H-F,H+F$ of odd type, using the same argument as in even type, then we finish the result. ∎
3.3. Embedding fake quadric of odd type in $\mathbb{P}^{4}$
In this subsection, we continue to study the fake quadrics in $\mathbb{P}^{4}$ and finish the remaining proof of the Theorem 1.4.
Proof of Theorem 1.4 of odd type..
The same notation as the even type, assume that $D\equiv xH+yF$ is an ample divisor. Therefore, the equation (2.3) implies that $x$ and $y$ must satisfy
$(x^{2}-y^{2})^{2}-10(x^{2}-y^{2})-5(3x+y)-4=0.$
It follows from Theorem 3.7 (ii) that $x>\rvert y\rvert$. By suitably modifying the previous equation, we can obtain that
$(x^{2}-y^{2}-5)^{2}=5(3x+y)+29,\quad\text{for}\quad x>\rvert y\rvert$ (3.5)
By [25, Chapter V, Corollary 2.9], we have that $h^{0}(S,\mathcal{O}_{S}(D))<\binom{4}{2}=6$. If $x+y>1$ and $x-y>2$, by Corollary 3.9 (i), we conclude that
$\frac{1}{2}(x+y-1)(x-y-2)=\chi(S,\mathcal{O}_{S}(D))\leq h^{0}(S,\mathcal{O}_{S}(D))<6$
unless $D\sim K_{S}$. This gives that
$(x+y-1)(x-y-2)\leq 10,\quad\text{for}\quad x+y>1,\quad x-y>2$ (3.6)
except for $x=3,y=-1$.
If $x+y\leq 1$ or $x-y\leq 2$ for $x>\rvert y\rvert$, there is no solution of equation (3.5). For $x+y>1$ and $x-y>2$, based on equations (3.5) and (3.6), it can be concluded that the values of $x,y$ are non-existent. Thus we finish the proof. ∎
Now, we turn our attention to the Corollary 1.5 of odd type.
Proof of Corollary 1.5 of odd type..
According to Theorem 3.7, the divisor $L\sim H$ is ample. Conisder the divisor $K_{S}+L$. Follows by Lemma 3.9, we have
$h^{0}(S,\mathcal{O}_{S}(K_{S}+L))=\chi(\mathcal{O}_{S}(K_{S}+L))=\chi(\mathcal{O}_{S}(4H-F))=\frac{1}{2}(4-1-1)(4+1-2)=3.$
If such divsior $K_{S}+L$ is very ample, then $\tau_{\rvert K_{S}+L\rvert}:S\to\mathbb{P}^{2}$ is an isomorphism, a contradiction.
Consider the divisor $2L$. By the exact sequence
$0\to h^{0}(S,\mathcal{O}_{S})\to h^{0}(S,\mathcal{O}_{S}(2L))\to h^{0}(2L,\mathcal{O}_{2L}(2L))\to 0,$
we only focus on $h^{0}(2L,\mathcal{O}_{2L}(2L))$. The Riemann-Roth Theorem gives that
$h^{0}(2L,\mathcal{O}_{2L}(2L))-h^{1}(2L,\mathcal{O}_{2L}(2L))=\deg(2L)\rvert_{2L}+1-p_{a}(2L)=4+1-6=-1.$
This shows that $h^{1}(2L,\mathcal{O}_{2L}(2L))\neq 0.$ Follows by Clifford’s Theorem
$h^{0}(2L,\mathcal{O}_{2L}(2L))-1=\dim|2L|_{2L}|<\frac{1}{2}\deg(2L)\rvert_{2L}=\frac{1}{2}\deg(2H)\rvert_{2H}=2,$
we get $h^{0}(2L,\mathcal{O}_{2L}(2L))<3.$ Using the above exact sequence, we have that
$h^{0}(S,\mathcal{O}_{S}(2L))=h^{0}(2L,\mathcal{O}_{2L}(2L))+1<4.$
If such divsior $2L$ is very ample, then $\tau_{\rvert 2L\rvert}:S\to\mathbb{P}^{2}$ is an isomorphism, a contradiction. ∎
3.4. The bounded cohomology property for fake quadrics of odd type
In this subsection, a curve also means a reduced and irreducible curve on a surface. Using the same argument as the even type (Sec.2.4), we can also prove the following result.
Proposition 3.10.
If $S$ is a fake quadric of odd type and a curve $C\equiv xH+yF,x,y\in\mathbb{Z}$, on $S$, then $C$ satisfies the bounded cohomology property for
$x+y>1\quad\text{and}\quad x-y>2,\quad\text{or}\quad x+y=1,\quad\text{or}\quad x-y=2,\quad\text{or}\quad x=\pm y.$
However, if $C$ satisfies $x=y+1$, then
$h^{0}(S,\mathcal{O}_{S}(C))-h^{1}(S,\mathcal{O}_{S}(C))=\chi(\mathcal{O}_{S}(C))=\frac{1}{2}(x+y-1)(x-y-2)=-y.$
This implies that the bounded cohomology property can hardly be determined for
$C\equiv(x+1)H+xF.$
Therefore we only consider that $S$ is a Mori dream space.
Proof of Theorem 1.6 of the odd type..
Under the above discussion, we only consider the curve $C\equiv(x+1)H+xF.$ Let $C_{1}\equiv H-F,\ C_{2}\equiv H+F$. If $S$ is a Mori dream space, by Corollary 1.2, we have that the linear system $|mC_{i}|$ is base point free for $m\gg 0.$ The $C^{2}=0$ gives that the Kodaira dimension $kod(S,C_{i})=1,i=1,2.$
Let $x_{0}$ be the minimal number such that $h^{0}(S,\mathcal{O}_{S}[(x+1)H+xF])>0$ for $x\geq x_{0}.$ Denote that
$C_{0}\equiv(x_{0}+1)H+x_{0}F\quad\text{and}\quad x=x_{0}+m,\quad m\gg 0.$
According to [14, 2.1.38], we have $h^{0}(S,\mathcal{O}_{S}(mC_{i}))\geq cm,$ where the const $c>0$ and $i=1,2$. Then we conclude
$h^{0}(S,\mathcal{O}_{S}(C))=h^{0}(S,\mathcal{O}_{S}(C_{0}+mC_{2}))\geq h^{0}(S,\mathcal{O}_{S}(mC_{2}))\geq cm$
and so $c^{-1}h^{0}(S,\mathcal{O}_{S}(C))\geq m.$ Recall that
$h^{0}(S,\mathcal{O}_{S}(C))+x=h^{1}(S,\mathcal{O}_{S}(C)).$
Thus we get that
$(c^{-1}+1)h^{0}(S,\mathcal{O}_{S}(C))\geq m+h^{0}(S,\mathcal{O}_{S}(C))=h^{1}(S,\mathcal{O}_{S}(C))-x_{0}.$
Since $h^{0}(S,\mathcal{O}_{S}(C))>x_{0}$ for $m$ sufficiently large, then we have
$(c^{-1}+2)h^{0}(S,\mathcal{O}_{S}(C))>(c^{-1}+1)h^{0}(S,\mathcal{O}_{S}(C))+x_{0}\geq h^{1}(S,\mathcal{O}_{S}(C)).$
This implies that the bounded cohomology property holds. ∎
Moreover, the above proof implies that
Corollary 3.11.
Let $S$ is a fake quadric of odd type. If the general fiber of the fibration $\pi:S\to\mathbb{P}^{1}$ is numerical equivalet to $m(H+F)$, $m>0,$ then the bounded cohomology property is true.